### v1

# Sums of four generalized polygonal numbers of almost prime length

Abstract
In this paper, we consider sums of four generalized $m$-gonal numbers whose parameters are restricted to integers with a bounded number of prime divisors. With some restriction on $m$ modulo 30, we show that for $n$ sufficiently large, it can be represented as such a sum, where the parameters are restricted to have at most 988 prime factors.
1 Introduction
For $m\in\mathbb{N}_{\geq 3}$ and $x\in\mathbb{Z}$, set
$p_{m}(x):=\frac{(m-2)x^{2}}{2}-\frac{(m-4)x}{2}.$
If $x\in\mathbb{N},$ then the number $p_{m}(x)$ counts the number of dots in a regular $m$-gon of length $x.$ In general, for $x\in\mathbb{Z},$ we call $p_{m}(x)$ the generalized $m$- gonal numbers.
For $n\in\mathbb{N},\alpha_{j}\in\mathbb{N},$ consider the equation
$\alpha_{1}p_{m}(x_{1})+\alpha_{2}p_{m}(x_{2})+\alpha_{3}p_{m}(x_{3})+\alpha_{4}p_{m}(x_{4})=n.$ (1.1)
It is obvious that if $(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4})$ is too arbitrary, i.e. $(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4})=(p,p,p,p).$ Then $n$ that can be represented is restricted. We hence restrict ourselves to the case where $\prod_{j=1}^{4}\alpha_{j}$ is odd and square-free. To understand the above sum and the relationship with shifted lattice theory, by completing the square, (1.1) is equivalent to
$\sum_{j=1}^{4}\alpha_{j}X_{j}^{2}=8(m-2)n+\sum_{j=1}^{4}\alpha_{j}(m-4)^{2},$
where $X_{j}=(2(m-2)x_{j}+4-m).$ Taking $L=2(m-2)\mathbb{Z}^{4}$ and $v=(4-m,...,4-m)\in\mathbb{Z}^{4},$ above is related to counting points of a given Euclidean distance on the shifted lattice $X=L+v.$
Theorem 1.1.
Suppose that $m$ is odd, $m-4\not\equiv 0\pmod{3}$ and $m-4\not\equiv 0\pmod{5}$. Further suppose $\text{ord}_{p}(\prod_{j=1}^{4}\alpha_{j})\leq 1$ for any odd prime $p.$ Then for $n$ sufficiently large, the equation (1.1) is solvable with each $x_{j}$ containing at most $988$ prime factors.
The paper is organized as follows. In Section 2, we recall the theory of quadratic lattices and shifted lattices and the interrelation between modular forms and theta series. In Section 3, we compute local densities for the Eisenstein series component of the theta series and obtain lower bounds for the Fourier coefficients of the Eisenstein series component. In section 4, we bound the Fourier coefficients coming from the contribution of the cuspidal part of the theta function. In Section 5, we obtain bounds to set up the sieve required in the proof of Theorem 1.1. We apply sieving techniques and prove Theorem 1.1 in Section 6.
2 Preliminaries
2.1 Setup and notation
Throughout this paper, we suppose $m$ to be odd. Let $V$ be a positive definite quadratic space over $\mathbb{Q}$ of rank $k$ associated with an non-degenerate symmetric bilinear form
$B:V\times V\rightarrow\mathbb{Q}\text{ with }Q(x)=B(x,x)\text{ for any }x\in V.$
A quadratic $\mathbb{Z}$-lattice is a free $\mathbb{Z}$-module
$L=\mathbb{Z}e_{1}+\mathbb{Z}e_{2}+\cdot\cdot\cdot+\mathbb{Z}e_{k}$
of rank $k,$ where $e_{1},...,e_{k}$ is a basis of $V.$ The dual lattice of $L$ is defined by
$L^{\#}:=\{v\in V:B(v,x)\in\mathbb{Z}\text{ }\forall x\in L\}.$
The $k\times k$ symmetric matrix $A_{L}:=(B(v_{i},v_{j}))$ is called the Gram matrix of $L$ with respect to a basis $v_{1},...,v_{k},$ and we write $L\cong(B(v_{i},v_{j})).$ We generally choose the vectors $v_{i}$ to be the standard basis element $e_{i}.$ The discriminant $d_{L}$ is defined to be the determinant of $(B(e_{i},e_{j})).$ If the Gram matrix is diagonal, then we write $L\cong\langle Q(e_{1}),...,Q(e_{k})\rangle.$ For a rational prime $p,$ we define the localization of $V$ and $L$ by
$V_{p}:=V\otimes_{\mathbb{Q}}\mathbb{Q}_{p}\text{ and }L_{p}:=L\otimes_{\mathbb{Z}}\mathbb{Z}_{p}.$
A lattice coset is a coset $L+v\in V/L$ where $v\in V.$ The conductor of the coset $L+v$ is the smallest integer $a$ such that $av\in L.$ Now consider
$f_{\boldsymbol{\alpha}}({\textbf{x}}):=\sum_{j=1}^{4}\alpha_{j}p_{m}(x_{j})$
and write
$h=h(n):=8(m-2)n+\sum_{j=1}^{4}\alpha_{j}(m-4)^{2}.$
Then $n$ is represented by $f$ iff $h$ is represented by $X^{\textbf{1}}=L^{\textbf{1}}+v$ where
$L^{\textbf{1}}=\mathbb{Z}(e_{1})+\mathbb{Z}(e_{2})+\mathbb{Z}(e_{3})+\mathbb{Z}(e_{4})\\ \cong\langle 4(m-2)^{2}\alpha_{1},4(m-2)^{2}\alpha_{2},4(m-2)^{2}\alpha_{3},4(m-2)^{2}\alpha_{4}\rangle$
and $v=\frac{4-m}{2(m-2)}(e_{1}+e_{2}+e_{3}+e_{4}).$ More generally, for ${\textbf{d}}=(d_{1},d_{2},d_{3},d_{4})\in\mathbb{N}^{4},$ define the coset $X^{\textbf{d}}:=L^{\textbf{d}}+v,$ where $L^{\textbf{d}}$ is the sublattice of $L^{\textbf{1}}$ defined by
$L^{\textbf{d}}=\mathbb{Z}(d_{1}e_{1})+\mathbb{Z}(d_{2}e_{2})+\mathbb{Z}(d_{3}e_{3})+\mathbb{Z}(d_{4}e_{4}).$
Then $n$ is represented by $f$ with the restriction $d_{j}\mid x_{j}$ iff $h$ is represented by $X^{\textbf{d}}.$ Therefore, we are interested in the number of solution to
$h=\sum_{j=1}^{4}(2(m-2)d_{j}x_{j}+4-m)^{2}.$
2.2 Modular forms
We will introduce modular form of half-integral weight in the following.
For positive integer N, we define
$\Gamma_{0}(N):=\{\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\Gamma_{0}(N):a,d\equiv 1\pmod{N}\}$,
$\Gamma_{1}(N):=\{\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\text{SL}_{2}(\mathbb{Z}):c\equiv 0\pmod{N}\}$,
to be two natural congruence subgroups of $\text{SL}_{2}(\mathbb{Z}).$
For $\gamma=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\Gamma_{0}(4)$ and $k\in\frac{1}{2}\mathbb{Z},$ we define the slash operator on a function $f:\mathbb{H}\rightarrow\mathbb{C}$ by
$f\mid_{k}\gamma(z)=(\frac{c}{d})^{2k}\epsilon_{d}^{2k}(cz+d)^{-k}f(\gamma z),$
where $\epsilon_{d}=\begin{cases}1,&\text{if $d\equiv 1\pmod{4}$}\\ i&\text{if $d\equiv 3\pmod{4}$}\end{cases}$, $(\frac{\cdot}{\cdot})$ is the Kronecker symbol, and $\gamma$ acts on $\mathbb{H}$ by fractional linear transformations defined by $\gamma z:=\frac{az+b}{cz+d}.$ Then we call $f$ a (holomorphic) modular form of weight $k$ on $\Gamma\subseteq\Gamma_{0}(4)$ where $\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\in\Gamma$ with character $\chi$ if
  1. 1.
    $f\mid_{k}\gamma=\chi(d)f$ for any $\gamma=\begin{pmatrix}a&b\\ c&d\end{pmatrix}$ in $\Gamma,$
  2. 2.
    $f$ is holomorphic on $\mathbb{H},$
  3. 3.
    $f(z)$ grows at most polynomially in $y$ as $z=x+iy\rightarrow\mathbb{Q}\cup\{i\infty\}.$
Then we call $f$ a cusp form if $f(z)\rightarrow 0$ as $z\rightarrow\mathbb{Q}\cup\{i\infty\}.$ Note that if $f$ is a modular form for a congruence group $\Gamma$ that contain $\begin{pmatrix}1&1\\ 0&1\end{pmatrix}$, then $f$ has a Fourier series expansion
$f(z)=\sum_{n=0}^{\infty}a(n)q^{n},$
where $q=e^{2\pi iz}.$ If $f$ is a cusp form, then $a(0)=0.$
2.3 Theta series for cosets
Let $X=L+v$ be a coset in a quadratic space $V$ of rank $k.$. For a positive integer $n$, we define
$R_{X}(n):=\{x\in X:Q(x)=n\},r_{X}(n):=\mid R_{X}(n)\mid.$
Then the theta series $\Theta_{X}(z)$ of the coset $X$ is defined as
$\Theta_{X}(z):=\sum_{x\in X}q^{Q(x)}=\sum_{n=0}^{\infty}r_{X}(n)q^{n}.$
Using [1, Proposition 2.1], a straightforward calculation shows that the theta series of cosets of rank $k$ are modular forms of weight $\frac{k}{2}.$ Therefore is splits into a sum of Eisenstein series $E_{X}$ and a cusp form $G_{X}.$ Let $a_{E_{X}}(n)$ and $a_{G_{X}}(n)$ denote the $n$-th Fourier coefficient of $E_{X}$ and $G_{X}$ for any integer $n\geq 0,$ respectively. We have
$r_{X}(n)=a_{E_{X}}(n)+a_{G_{X}}(n).$
3 Eisenstein series components
Let $X=L+v$ be a lattice coset of rank $k,$ and let $h\in\mathbb{Q}.$ For $p$ prime and $z\in\mathbb{Q}_{p},$ we define $\textbf{e}_{p}(z)=e^{2\pi i(-y)}$ with $y\in\cup_{t=1}^{\infty}p^{-t}\mathbb{Z}$ such that $z-y\in\mathbb{Z}_{p}.$ Let $\lambda(v)$ and $\lambda_{p}$ be the characteristic functions of $X$ and $X_{p}$ respectively. Normalize the measures $dv$ and $d\sigma$ on $L_{p}$ and $\mathbb{Z}_{p}$ so that $\int_{L_{p}}\,dv=\int_{\mathbb{Z}_{p}}\,dv=1.$ Define the local density at p by
$b_{p}(h,\lambda,0):=\int_{\mathbb{Q}_{p}}\int_{L_{p}}\textbf{e}_{p}(\sigma(Q(v+\nu)-h)\,dv\,d\sigma.$ (3.1)
Shimura [7, Theorem 1.5] showed that the $h$-th Fourier coefficient of the Eisenstein series part of $\Theta_{X}$ can be expressed as the product of local densities. We refer readers to [8, 9] for more details.
We now provide the statement for the special case when $k=4.$
Theorem 3.1.
Let $X=L+v$ be a quaternary lattice coset, and let $h\in\mathbb{Q}.$ Then
$a_{E_{X}}(h)=\frac{(2\pi)^{2}h}{[\frac{1}{2}L^{\#}:L]^{\frac{1}{2}}\Gamma(2)L(2,\psi)}\cdot\prod_{p\mid e_{1}}\frac{b_{p}(h,\lambda,0)}{(1-\psi(p)p^{-2})}\cdot\prod_{{p\mid h\atop{p\nmid e_{1}}}}\gamma_{p}(2).$
Here $\psi(\cdot)$ is primitive character associated to the real character $(\frac{-1}{\cdot}),$ $e_{1}$ is the product of all primes p at which $h\not\in\mathbb{Z}_{p},$ $e^{\prime}$ is the product of all primes such that $L^{\#}\neq 2L_{p},$ the numbers $\gamma_{p}(s)$ are given in [7, Section 1.6]. Let $\lambda_{\textbf{d}}$ be the characteristic function of $X^{\textbf{d}}.$ In our case $e_{1}$ is the product of all primes dividing $2(m-2)\prod_{j=1}^{4}\alpha_{j}d_{j},$ and $e_{1}=e^{\prime}.$ Therefore we have
$a_{E_{X}}(h)=\frac{(2\pi)^{2}h}{(16d_{L^{\textbf{d}}})^{\frac{1}{2}}\Gamma(2)L(2,\psi)}\cdot\prod_{p\mid e_{1}}\frac{b_{p}(h,\lambda,0)}{(1-\psi(p)p^{-2})}\cdot\prod_{{p\mid h\atop{p\nmid e_{1}}}}\gamma_{p}(2).$ (3.2)
3.1 Computation of local densities of $X^{\textbf{d}}$.
In this subsection, we compute the local densities $b_{p}(h,\lambda_{\textbf{d}},0)$ of $X^{\textbf{d}}$ in order to evaluate $a_{E_{X^{\textbf{d}}}}(h)$ using (3.2).
We first compute the case for $p=2.$
Lemma 3.2.
The local density if $X^{\textbf{d}}$ at the prime 2 is given by
$b_{p}(h,\lambda_{\textbf{d}},0)=\begin{cases}2^{2+\text{min}_{j}\{\text{ord}_{2}(\alpha_{j}d_{j})\}},&\text{if $8(m-2)n\in 4\gcd(\alpha_{1}d_{1},\alpha_{2}d_{2},\alpha_{3}d_{3},\alpha_{4}d_{4})\mathbb{Z}_{2}$,}\\ 0,&\text{otherwise.}\end{cases}$
Proof.
By [1 Lemma 3.3]. ∎
For odd $p\mid(m-2)$
Lemma 3.3.
The local density if $X^{\textbf{d}}$ at odd prime $p\mid(m-2)$ is given by
$b_{p}(h,\lambda_{\textbf{d}},0)=\begin{cases}p^{\text{ord}_{p}(m-2)+\text{min}\{\text{ord}_{p}(\alpha_{j}d_{j})\}},&\text{if $n\in\gcd(\alpha_{1}d_{1},\alpha_{2}d_{2},\alpha_{3}d_{3},\alpha_{4}d_{4})\mathbb{Z}_{p}$,}\\ 0,&\text{otherwise.}\end{cases}$
Proof.
By [1 Lemma 3.4]. ∎
For $p\geq 3,$ we will use the formula from [5, Theorem 4.2] to compute the local densities. We provide the statement for $r=4$ in the following.
Theorem 3.4.
Let $p\geq 3$ be an odd prime. Suppose that $n\in\mathbb{Z}_{p}$ and $\phi(\textbf{x})=\sum_{i=1}^{4}(b_{i}x_{i}^{2}+c_{i}x_{i})$ with $b_{i},c_{i}\in\mathbb{Z}_{p}$ for $1\leq i\leq 4.$ For $1\leq i\leq 4,$ we define $t_{i}=\text{min}(\text{ord}_{p}(b_{i}),\text{ord}_{p}(c_{i})).$ We set
$D_{p}:=\{1\leq i\leq 4\mid\text{ord}_{p}(b_{i})>\text{ord}_{p}(c_{i})\},$
$N_{p}:=\{1\leq i\leq 4\mid\text{ord}_{p}(b_{i})\leq\text{ord}_{p}(c_{i})\}.$
and set $\textbf{t}_{d}:=\text{min}\{t_{i}\mid i\in D_{p}\}.$ We further define
$\textbf{n}:=n+\sum_{i\in N_{p}}\frac{c_{i}^{2}}{4b_{i}}.$
If $\textbf{n}\neq 0,$ we assume that $\textbf{n}=\textbf{u}_{n}p^{\textbf{t}_{n}}$ with $\textbf{u}\in\mathbb{Z}_{p}^{\times}$ and $\textbf{t}_{n}\in\mathbb{Z}.$ Otherwise we set $\textbf{t}_{n}:=\infty.$ For an integer $t\in\mathbb{Z},$ we define
$L_{p}(t):=\{i\in N_{p}\mid t_{i}-t<0\text{ and odd}\},l_{p}(t):=\mid L_{p}(t)\mid.$
Then we have
$I_{p}(n;\phi)=1+(1-\frac{1}{p})\sum_{1\leq t\leq\text{min}(\textbf{t}_{d},\textbf{t}_{n})\atop l_{p}(t)\text{ even}}\delta_{p}(t)p^{\tau_{p}(t)}+\delta_{p}(\textbf{t}_{n}+1)w_{p}p^{\tau_{p}(\textbf{t}_{n}+1)},$
where for any integer $t\in\mathbb{Z},$ we define
$\delta_{p}(t):=\epsilon_{p}^{3l_{p}(t)}\prod_{i\in L_{p}(t)}(\frac{u_{i}}{p}),\tau_{p}(t):=t+\sum_{i\in N_{p}\atop t_{i}<t}\frac{t_{i}-t}{2},$
with $u_{i}:=p^{-\text{ord}(b_{i})}b_{i}\in\mathbb{Z}_{p}^{\times}$ and define
[…]
Using Theorem 3.4, we have the following result.
Lemma 3.5.
Let $p$ be an odd prime such that $p\nmid(m-2)(m-4),p\mid\prod_{j=1}^{4}d_{j},$ and $p\nmid\prod_{j=1}^{4}\alpha_{j}$
When $\mid N_{p}\mid=0,$ we have
$b_{p}(h,\lambda_{\textbf{d}},0)=\begin{cases}p&\text{if $N\in p\mathbb{Z}_{p}$,}\\ 0&\text{if $N\not\in p\mathbb{Z}_{p}$.}\end{cases}$
When $\mid N_{p}\mid=1,$ we have
$0\leq b_{p}(h,\lambda_{\textbf{d}},0)\leq 2.$
When $\mid N_{p}\mid=2,$ we have
$\frac{1}{p}\leq b_{p}(h,\lambda_{\textbf{d}},0)\leq 2-\frac{1}{p}.$
When $\mid N_{p}\mid=3,$ we have
$1-\frac{1}{p}\leq b_{p}(h,\lambda_{\textbf{d}},0)\leq 1+\frac{1}{p}.$
Lower bound for the $\textbf{d}=1$ case. We are now ready to give a lower bound for $a_{E_{X^{\textbf{1}}}(h)}.$
Lemma 3.6.
We have
$a_{E_{X^{\textbf{1}}}}(h)\gg_{m,\epsilon}h^{1-\epsilon}.$
Proof.
By using (3.2) to the case $\textbf{d}=1,$ we have
$a_{E_{X^{\textbf{1}}}}(h)=\frac{(2\pi)^{2}h}{(16d_{L^{\textbf{1}}})^{\frac{1}{2}}\Gamma(2)L(2,\psi)}\cdot\prod_{p\mid 2(m-2)\prod_{j=1}^{4}\alpha_{j}}\frac{b_{p}(h,\lambda,0)}{(1-\psi(p)p^{-2})}\cdot\prod_{{p\mid h\atop{p\nmid 2(m-2)\prod_{j=1}^{4}\alpha_{j}}}}\gamma_{p}(2).$
By using Lemma 3.3 and Lemma 3.4, we have
$\prod_{p\mid 2(m-2)\prod_{j=1}^{4}\alpha_{j}}\frac{b_{p}(h,\lambda,0)}{(1-\psi(p)p^{-2})}$ $=\frac{16(m-2)}{3}\prod_{p\mid(m-2)\prod_{j=1}^{4}\alpha_{j}}\frac{p^{\text{min}_{j}\{\text{ord}_{p}(\alpha_{j})\}}}{1-\psi(p)p^{-2}}$
$\geq\frac{16(m-2)}{3}\prod_{p\mid(m-2)\prod_{j=1}^{4}\alpha_{j}}\frac{1}{1+p^{-2}}$
For $p\mid h$ and $p\nmid 2(m-2)\prod_{j=1}^{4}\alpha_{j},$ by [6, (3.30)] we have the following bound
$\prod_{{p\mid h\atop{p\nmid 2(m-2)\prod_{j=1}^{4}\alpha_{j}}}}\gamma_{p}(2)\geq\prod_{p\leq h}(1-\frac{1}{p})\gg h^{-\epsilon}$
Hence we have
$a_{E_{X^{\textbf{1}}}}(h)\gg_{m,\epsilon}h^{1-\epsilon}.$
3.2 Ratio of Eisenstein series.
Let $S_{2^{j}}$ be the set of $d\in\mathbb{N}$ with either $d$ odd and squarefree or $d=2^{j}d^{\prime}$ with $d^{\prime}$ odd and squarefree. From (3.2), we have for $\textbf{d}\in S_{1}^{4}$ and $\boldsymbol{\ell}\in S_{1}^{4}$ with $\gcd(d_{1}d_{2}d_{3}d_{4},\ell_{1}\ell_{2}\ell_{3}\ell_{4})=1$ that
$\begin{split}&\frac{a_{E_{X^{\boldsymbol{d\ell}}}(h)}}{a_{E_{X^{\boldsymbol{\ell}}}(h)}}\\ &=\frac{1}{d_{1}d_{2}d_{3}d_{4}}\cdot\prod_{p\mid(m-2)}\frac{b_{p}(h,\lambda_{\boldsymbol{d\ell}},0)}{b_{p}(h,\lambda_{\boldsymbol{\ell}},0)}\cdot\prod_{p\mid d_{1}d_{2}d_{3}d_{4}\atop p\nmid 2(m-2)}\frac{b_{p}(h,\lambda_{\textbf{d}},0)}{1-\psi(p)p^{-2}}\cdot\prod_{p\mid(h,d_{1}d_{2}d_{3}d_{4})\atop p\nmid 2(m-2)}\gamma_{p}^{-1}(2).\end{split}$ (3.3)
Define
$\begin{split}\beta_{X^{\textbf{p}^{c}},p}(h):=\frac{1}{p^{c_{1}+c_{2}+c_{3}+c_{4}}}&(\frac{b_{p}(h,\lambda_{\textbf{p}^{c}},0)}{b_{p}(h,\lambda_{\textbf{1}},0)})^{\delta_{p\mid(m-2)}}\\ &\cdot(b_{p}(h,\lambda_{\textbf{p}^{c}},0)\cdot\frac{1}{1-\psi(p)p^{-2}}\cdot\gamma_{p}^{-1}(2))^{\delta_{p\nmid 2(m-2),c\neq 0}}\end{split}$ (3.4)
Therefore we have
$\frac{a_{E_{X^{\boldsymbol{d\ell}}}(h)}}{a_{E_{X^{\boldsymbol{\ell}}}(h)}}=\prod_{p\text{ odd}}\beta_{X^{\textbf{p}^{ord_{p}(\textbf{d})}},p}(h),$ (3.5)
where the product runs over all odd prime numbers.
Lemma 3.7.
Let $m$ be an odd integer such that $m\not\equiv 4\pmod{3},m\not\equiv 4\pmod{5}$, $p$ be an odd prime, $h=8(m-2)n+\sum_{j=1}^{4}\alpha_{j}(m-4)^{2}.$ Then for $p\nmid(m-2)(m-4)$ and $p\nmid\prod_{j=1}^{4}\alpha_{j}$ we have the following:
$\beta_{X^{(p,1,1,1),p}}(h)\leq\frac{2}{p}$
$\beta_{X^{(p,p,1,1),p}}(h)\leq\frac{4}{p^{2}}$
$\beta_{X^{(p,p,p,1),p}}(h)\leq\frac{8}{p^{3}}$
$\beta_{X^{(p,p,p,p),p}}(h)\leq\begin{cases}\frac{1}{p^{2}(p-1)},&\text{for $n\in p\mathbb{Z}_{p}$ }\\ \frac{16}{p^{4}},&\text{for $n\not\in p\mathbb{Z}_{p}$ }\end{cases}$
Then we have
$\prod_{p\mid\textbf{d}}\beta_{X^{\textbf{d}},p}(h)\leq\frac{\tilde{w}(d_{1})\tilde{w}(d_{2})\tilde{w}(d_{3})\tilde{w}(d_{4})}{d_{1}d_{2}d_{3}d_{4}}$
where $\frac{\tilde{w}(p)}{p}:=\begin{cases}\frac{2}{p},&\text{if $p\nmid n$,}\\ \max\{(\frac{1}{p^{2}(p-1)})^{\frac{1}{4}},\frac{2}{p}\},&\text{if $p\mid n$}\end{cases}$
Proof.
Let
$w_{p}(h):=\frac{1}{1-\psi(p)p^{-2}}\gamma_{p}(2)^{-1}$
Then by [7] we have
$\gamma_{p}(2)\geq 1-\frac{1}{p}.$
Thus
$w_{p}(h)\leq\frac{1}{1-p^{-2}}\frac{p}{p-1}=\frac{p^{3}}{(p-1)^{2}(p+1)}\text{ for }p\geq 5$
and
[…]
Therefore we have
$\beta_{X^{(p,1,1,1),p}}(h)$ $=\frac{1}{p}w_{p}(h)b_{p}(h,\lambda_{(p,1,1,1)},0)$
$\leq\frac{2}{p}$
and similarly for other cases. ∎
Lemma 3.8.
Let $g(\textbf{d}):=\prod_{p\mid\textbf{d}}\frac{\beta_{X^{\textbf{d}},p}(h)}{\prod_{j=1}^{4}\beta_{X^{(d_{j},1,1,1)},p}(h)},u_{i,j}:=\gcd(d_{i},d_{j}).$ Then we have
$g(\textbf{d})\ll\max_{i,j}(u_{i,j})^{12}.$
Proof.
By Lemma 3.7, we have
$\beta_{X^{\textbf{d}},p}(h)\leq\begin{cases}(\frac{2}{p})^{\sum_{i=1}^{4}{\text{ord}_{p}(d_{i})}},&\text{if $p\nmid n$ }\\ (\frac{1}{\sqrt{p}})^{\sum_{i=1}^{4}{\text{ord}_{p}(d_{i})}},&\text{if $p\mid n$}\end{cases}$
and we have
$\beta_{X^{(p,1,1,1)},p}(h)\geq\frac{p-1}{2p^{2}-p+1}.$
Hence we have
$g(\textbf{d})$ $=\prod_{p\mid u_{1,2}\cdot\cdot\cdot u_{3,4}}\frac{\beta_{X^{\textbf{d}},p}(h)}{\prod_{j=1}^{4}\beta_{X^{(p,1,1,1)},p}(h)}$
$\leq\prod_{p\mid u_{1,2}\cdot\cdot\cdot u_{3,4}}(\frac{2p^{2}-p+1}{\sqrt{p}(p-1)})^{\text{ord}_{p}(d_{1})+\text{ord}_{p}(d_{2})+\text{ord}_{p}(d_{3})+\text{ord}_{p}(d_{4})}$
$\leq\prod_{p\mid u_{1,2}\cdot\cdot\cdot u_{3,4}}(\frac{2p^{2}-p+1}{\sqrt{p}(p-1)})^{4}$
$\leq\prod_{p\mid u_{1,2}\cdot\cdot\cdot u_{3,4}}(\frac{4^{4}(p-1)^{8}}{p^{2}(p-1)^{4}})$
$\leq\prod_{p\mid u_{1,2}\cdot\cdot\cdot u_{3,4}}4^{4}p^{2}\leq 4^{4}\prod_{i<j\atop 1\leq i,j\leq 4}u_{i,j}^{2}\ll\max_{i,j}(u_{i,j})^{12}$
Lemma 3.9.
Let $m$ be an odd integer such that $m\not\equiv 4\pmod{3},m\not\equiv 4\pmod{5}$, $p\geq 5$ be an odd prime, $h=8(m-2)n+\sum_{j=1}^{4}\alpha_{j}(m-4)^{2}.$ Then for $p\nmid(m-2)(m-4)$ and $p\mid\prod_{j=1}^{4}\alpha_{j}$ we have the following:
$\max\{\beta_{X^{(p,1,1,1)},p}(h),\beta_{X^{(1,p,1,1)},p}(h),\beta_{X^{(1,1,p,1)},p}(h),\beta_{X^{(1,1,1,p)},p}(h)\}\leq\frac{4}{p}$
$\max\{\beta_{X^{(p,p,1,1)},p}(h),\beta_{X^{(p,1,p,1)},p}(h),\beta_{X^{(p,1,1,p)},p}(h),\beta_{X^{(1,p,p,1)},p}(h)\\ ,\beta_{X^{(1,p,1,p)},p}(h),\beta_{X^{(1,1,p,p)},p}(h)\}\leq\frac{2p}{(p-1)^{2}(p+1)}\leq\frac{16}{p^{2}}$
$\max\{\beta_{X^{(p,p,p,1)},p}(h),\beta_{X^{(p,p,1,p)},p}(h),\beta_{X^{(p,1,p,p)},p}(h),\beta_{X^{(1,p,p,p)},p}(h)\}\leq\frac{2}{(p-1)^{2}(p+1)}\leq\frac{64}{p^{3}}$
$\beta_{X^{(p,p,p,p)},p}(h)\leq\begin{cases}\frac{1}{(p-1)^{2}(p+1)},&\text{for $n\in p\mathbb{Z}_{p}$ }\\ \frac{256}{p^{4}},&\text{for $n\not\in p\mathbb{Z}_{p}$ }\end{cases}$
Proof.
By Theorem 3.4, we have
$\max\{\beta_{X^{(p,1,1,1)},p}(h),\beta_{X^{(1,p,1,1)},p}(h),\beta_{X^{(1,1,p,1)},p}(h),\beta_{X^{(1,1,1,p)},p}(h)\}\\ \leq\frac{1}{p}\cdot(2-\frac{1}{p})\cdot\frac{p^{3}}{(p-1)^{2}(p+1)}\leq\frac{4}{p}\text{ for }p\geq 5.$
and similarly for other cases. ∎
4 Error from Cusp form
We will use the formula from [4, Lemma 3.2]. We provide the statement for the case $l=4.$
Lemma 4.1.
Let $M=2(m-2),N_{\boldsymbol{\alpha}}$ be the level of $f_{\boldsymbol{\alpha}}(\textbf{x}).$ Then for any $\delta,\epsilon>0,$ there exists constants $c_{\delta},C_{\epsilon}>0$ such that
$\mid a_{G_{X}}(n)\mid\leq\frac{54}{\pi^{2}\delta_{M^{2}N_{\boldsymbol{\alpha}}}^{\frac{3}{2}}}\frac{M^{2}N_{\boldsymbol{\alpha}}^{2+2\delta}\sqrt{\frac{2\pi}{3}}e^{2\pi}\zeta(1+4\delta)^{\frac{1}{2}}c_{\delta}^{\frac{5}{2}}\varphi(M)}{\prod_{p\mid M,p\nmid N_{\boldsymbol{\alpha}}}(1-p^{-2})^{\frac{1}{2}}\prod_{p\mid N_{\boldsymbol{\alpha}}}(1-p^{-1})^{\frac{1}{2}}}C_{\epsilon}n^{\frac{1}{2}+\epsilon}\\ \times(\sum_{d\mid M^{2}N_{\boldsymbol{\alpha}}}\varphi(\frac{M^{2}N_{\boldsymbol{\alpha}}}{d})\varphi(d)\frac{M^{2}N_{\boldsymbol{\alpha}}}{d}(\frac{\gcd(M^{2},d)}{M^{2}})^{4})^{\frac{1}{2}}(\frac{27}{\Delta_{\boldsymbol{\alpha}}}\frac{M^{2}N_{\boldsymbol{\alpha}}}{\pi\delta_{M^{2}N_{\boldsymbol{\alpha}}}}+16)^{\frac{1}{2}}.$
Using Lemma 4.1, we have the following bound.
Lemma 4.2.
$\mid a_{G_{X}}(n)\mid\ll M^{\frac{11}{2}+\epsilon}N_{\boldsymbol{\alpha}}^{\frac{5}{2}+\epsilon}n^{\frac{1}{2}+\epsilon}.$
Proof.
Note that we have
$\sum_{d\mid M^{2}N_{\boldsymbol{\alpha}}}\varphi(\frac{M^{2}N_{\boldsymbol{\alpha}}}{d})\varphi(d)\frac{M^{2}N_{\boldsymbol{\alpha}}}{d}(\frac{\gcd(M^{2},d)}{M^{2}})^{4}$ $\leq\sum_{d\mid M^{2}N_{\boldsymbol{\alpha}}}\varphi(\frac{M^{2}N_{\boldsymbol{\alpha}}}{d})\varphi(d)\frac{M^{2}N_{\boldsymbol{\alpha}}}{d}$
$\leq M^{2}N_{\boldsymbol{\alpha}}\varphi(M^{2}N_{\boldsymbol{\alpha}})\sigma_{-1}(M^{2}N_{\boldsymbol{\alpha}})\ll_{\epsilon}M^{4+\epsilon}$
Therefore we have
$\mid a_{G_{X}}(n)\mid$ $\ll M^{\frac{5}{2}}N_{\boldsymbol{\alpha}}^{2+\epsilon}\log(\log(N_{\boldsymbol{\alpha}}))\log(\log(M))\cdot M^{2+\epsilon}\cdot MN_{\tilde{a}}^{\frac{1}{2}}n^{\frac{1}{2}+\epsilon}$
$\ll M^{\frac{11}{2}+\epsilon}N_{\boldsymbol{\alpha}}^{\frac{5}{2}+\epsilon}n^{\frac{1}{2}+\epsilon}$
5 Main term and error term from sieving
Define $A_{\boldsymbol{\ell}}$ be the set of solutions ${\textbf{x}}\in\mathbb{Z}^{4}$ with $\ell_{j}\mid x_{j}$ to
$\sum_{j=1}^{4}\alpha_{j}p_{m}(x_{j})=n.$
Let $P$ be a finite set of odd primes and for $\boldsymbol{\ell}\in\mathbb{N}^{3}$ and $n\in\mathbb{N},$ let $\mathcal{S}_{h}({A_{\boldsymbol{\ell}}},P)$ denote the size of the set
$\mathcal{F}_{h}({A_{\boldsymbol{\ell}}},P)=\#\{\textbf{x}\in\mathbb{Z}^{4}:\sum_{j=1}^{4}\alpha_{j}p_{m}(\ell_{j}x_{j})=n,p\mid x_{j}\implies p\not\in P\}.$ (5.1)
We then define
$\mathcal{S}_{{\textbf{c}},h}({A_{\boldsymbol{\ell}}},P):=\sum_{p\mid{2\prod_{j=1}^{4}\alpha_{j}}}\sum_{\textbf{b}(p)\in\{0,1\}^{4}}\prod_{i=1}^{4}\mu(p^{b(p)_{i}})\mathcal{S}_{h}(A_{{\textbf{t}}\cdot\boldsymbol{\ell}},P),$ (5.2)
where ${\textbf{t}}=\prod_{p\mid 2\prod_{j=1}^{4}\alpha_{j}}p^{c_{p}\cdot{\textbf{b}(p)}}.$
Which is the size of the set
$\mathcal{F}_{c,h}({A_{\boldsymbol{\ell}}},P)=\{\textbf{x}\in\mathbb{Z}^{4}:\sum_{j=1}^{4}\alpha_{j}p_{m}(\ell_{j}x_{j})=n,\\ p\mid x_{j}\implies p\not\in P\cup\{p:p\mid 2\prod_{j=1}^{4}\alpha_{j}\}\\ \mbox{ or }(p\mid 2\prod_{j=1}^{4}\alpha_{j}\mbox{ and }\text{ord}_{p}(x_{j})<c_{p})\}.$
Define $P=P_{z,\Delta_{\alpha}}:=\{p\leq z\}\cap\{p\nmid 2\prod_{j=1}^{4}\alpha_{j}\}.$
For $z_{0}\geq 3,$ define $P(z_{0}):=\prod_{p<z_{0}\atop p\nmid 2\prod\alpha_{j}}p$
For $\beta,D>0$ and the integer $d$ of the form $d=p_{1}\cdot\cdot\cdot p_{r}$ with $p_{1}>\cdot\cdot\cdot>p_{r}$ with $p_{j}$ an odd prime. The Rosser weights $\lambda_{d}^{\pm}$ are defined as follows:
Let
$y_{m}=y_{m}(D,\beta):=(\frac{D}{p_{1}\cdot\cdot\cdot p_{m}})^{\frac{1}{\beta}},$
then
$\lambda_{d}^{+}=\lambda_{d,D}^{+}(\beta):=\begin{cases}(-1)^{r}\quad\text{if $p_{2l+1}<y_{2l+1}(D,\beta)\quad\forall 0\leq l\leq\frac{r-1}{2}$},\\ 0\quad\qquad\text{otherwise.}\end{cases}$
$\lambda_{d}^{-}=\lambda_{d,D}^{-}(\beta):=\begin{cases}(-1)^{r}\quad\text{if $p_{2l}<y_{2l}(D,\beta)\quad\forall 0\leq l\leq\frac{r}{2}$},\\ 0\quad\qquad\text{otherwise.}\end{cases}$
Furthermore define
$\Lambda_{d}^{-}:=4\lambda_{d}^{-}-3\lambda_{d}^{+}.$
By [2], we have the following inequalities
$\sum_{d\mid c}\lambda_{d}^{-}\leq\sum_{d\mid c}\mu(d)\leq\sum_{d\mid c}\lambda_{d}^{+}$ (5.3)
and
$\prod_{j=1}^{4}\sum_{d_{j}\mid c_{j}}\mu(d_{j})\geq\sum_{k=1}^{4}\sum_{d_{k}\mid c_{k}}\lambda_{d_{k}}^{-}\prod_{1\leq 4\atop{j\neq k}}\sum_{d_{j}\mid c_{j}}\lambda_{d_{j}}^{+}-3\prod_{j=1}^{4}\sum_{d_{j}\mid c_{j}}\lambda_{d_{j}}^{+}.$ (5.4)
Define $W_{\textbf{c},h}(z_{0}):=\prod_{p\mid 2\prod_{j=1}^{4}\alpha_{j}}C_{p}\prod_{p\in P_{z_{0},\Delta_{\alpha}}}(1-\beta_{X^{(p,1,1,1)},p}(h))$ and
$H(n)=\prod_{p\mid n}(1+p^{-\frac{1}{2}}).$
where $C_{p}$ are the ratio of local densities at $p.$ For $\epsilon\in\{\pm 1\},$ we set
$M_{h}^{\epsilon}(z_{0})=\begin{cases}\sum_{d_{1}\mid P_{z_{0},\Delta_{\alpha}}}\Lambda_{d_{1}}^{-}\sum_{d_{2}\mid P_{z_{0},\Delta_{\alpha}}}\lambda_{d_{2}}^{+}\sum_{d_{3}\mid P_{z_{0},\Delta_{\alpha}}}\lambda_{d_{3}}^{+}\\ \qquad\qquad\qquad\qquad\qquad\cdot\sum_{d_{4}\mid P_{z_{0},\Delta_{\alpha}}}\lambda_{d_{4}}^{+}\prod_{p\mid\textbf{d}}\beta_{X^{\textbf{d}},p}(h)&\text{if $\epsilon=-1$},\\ \sum_{d_{1}\mid P_{z_{0},\Delta_{\alpha}}}\lambda_{d_{1}}^{+}\sum_{d_{2}\mid P_{z_{0},\Delta_{\alpha}}}\lambda_{d_{2}}^{+}\sum_{d_{3}\mid P_{z_{0},\Delta_{\alpha}}}\lambda_{d_{3}}^{+}\\ \qquad\qquad\qquad\qquad\qquad\cdot\sum_{d_{4}\mid P_{z_{0},\Delta_{\alpha}}}\lambda_{d_{4}}^{+}\prod_{p\mid\textbf{d}}\beta_{X^{\textbf{d}},p}(h)&\text{if $\epsilon=1$},\end{cases}$
Proposition 5.1.
[…]
$S_{\textbf{c},h}(A_{\boldsymbol{\ell}},P_{z_{0},\Delta_{\alpha}}))=(W_{c,h}(z_{0})+$ $O_{\epsilon}(\Delta^{-\frac{1}{2}}H(n)^{5}\log(z_{0})^{8}$
$+\Delta^{\epsilon-B}\log(z_{0})^{8}+\Delta^{C}e^{-s_{0}})r_{gen^{+}(X^{\boldsymbol{\ell}})}(h)$
$+O(\sum_{{\textbf{d}\in S_{\prod_{p\mid 2\prod_{j=1}^{4}\alpha_{j}}p^{c_{p}}}^{4}\atop{p\mid d_{j}\implies p\in P_{\delta,\Delta_{\alpha}}}}}\mid r_{X^{{\textbf{d}}\cdot{\boldsymbol{\ell}}}}(h)-r_{gen^{+}(X^{{\textbf{d}}\cdot{\boldsymbol{\ell}}})}(h)\mid).$
Before proving Proposition 5.1, we require certain inequalities for the upper bound and the lower bound of $S_{\textbf{c},h}(A_{\ell},P_{z_{0},\Delta_{\alpha}}).$
Proposition 5.2.
[…]
(1) We have
$S_{\textbf{c},h}(A_{\ell},P_{z_{0},\Delta_{\alpha}})\geq\prod_{p\mid 2\prod_{j=1}^{4}\alpha_{j}}C_{p}M_{h}^{-}(n^{\delta})r_{gen^{+}(X^{\boldsymbol{\ell}})}(h)\\ -7\sum_{{\textbf{d}\in S_{\prod_{p\mid 2\prod_{j=1}^{4}\alpha_{j}}p^{c_{p}}}^{4}\atop{p\mid d_{j}\implies p\in P_{\delta,\Delta_{\alpha}}}}}\mid(r_{X^{\textbf{d}\cdot\boldsymbol{\ell}}}(h)-r_{gen^{+}(X^{\textbf{d}\cdot\boldsymbol{\ell}})}(h))\mid.$
(2) We have
$S_{\textbf{c},h}(A_{\ell},P_{z_{0},\Delta_{\alpha}})\leq\prod_{p\mid 2\prod_{j=1}^{4}\alpha_{j}}C_{p}M_{h}^{+}(n^{\delta})r_{gen^{+}(X^{\boldsymbol{\ell}})}(h)\\ +7\sum_{{\textbf{d}\in S_{\prod_{p\mid 2\prod_{j=1}^{4}\alpha_{j}}p^{c_{p}}}^{4}\atop{p\mid d_{j}\implies p\in P_{\delta,\Delta_{\alpha}}}}}\mid(r_{X^{\textbf{d}\cdot\boldsymbol{\ell}}}(h)-r_{gen^{+}(X^{\textbf{d}\cdot\boldsymbol{\ell}})}(h))\mid.$
Proof.
(1) Suppose $P_{1}$ and $P_{2}$ are two disjoint sets of primes with $p\nmid l_{j}$ for all $p\in P_{2}$ and $1\leq j\leq 4.$ Writing $R=\prod_{p\in P_{2}}p,$ we compute
$\mathcal{S}_{\textbf{c},h}(A_{\ell},P_{1}\cup P_{2})$ $=\sum_{x\in F_{\textbf{c},h}(A_{\ell},P_{1})}\prod_{j=1}^{4}\sum_{d_{j}\mid\gcd(x_{j},R)}\mu(d_{j})$
$=\sum_{x\in F_{\textbf{c},h}(A_{\ell},P_{1})}\sum_{{d_{1}\mid R\atop{d_{1}\mid x_{1}}}}\mu(d_{1})\sum_{{d_{2}\mid R\atop{d_{2}\mid x_{2}}}}\mu(d_{2})\sum_{{d_{3}\mid R\atop{d_{3}\mid x_{3}}}}\mu(d_{3})\sum_{{d_{4}\mid R\atop{d_{4}\mid x_{4}}}}\mu(d_{4})$
$\geq\sum_{x\in F_{\textbf{c},h}(A_{\ell},P_{1})}\sum_{{d_{1}\mid R\atop{d_{1}\mid x_{1}}}}\Lambda^{-}_{d_{1}}\sum_{{d_{2}\mid R\atop{d_{2}\mid x_{2}}}}\lambda^{+}_{d_{2}}\sum_{{d_{3}\mid R\atop{d_{3}\mid x_{3}}}}\lambda^{+}_{d_{3}}\sum_{{d_{4}\mid R\atop{d_{4}\mid x_{4}}}}\lambda^{+}_{d_{4}}$
$=\sum_{d_{1}\mid R}\Lambda^{-}_{d_{1}}\sum_{d_{2}\mid R}\lambda^{+}_{d_{2}}\sum_{d_{3}\mid R}\lambda^{+}_{d_{3}}\sum_{d_{4}\mid R}\lambda^{+}_{d_{4}}\mathcal{S}_{{\textbf{c}},h}(A_{\textbf{d}\cdot\boldsymbol{\ell}},P_{1}),$
using (5.4) in the second last step. We assume that $p\mid\ell_{j}\implies p>z_{0}$ and $p\mid\prod_{j=1}^{4}\alpha_{j}$ and take $P_{1}=\emptyset,P_{2}=P_{z_{0},\Delta_{\alpha}}:=\{p\leq z_{0},p\nmid 2\prod_{j=1}^{4}\alpha_{j}\},$ then by (5.2) we have
$S_{\textbf{c},h}(A_{{\textbf{d}\cdot\boldsymbol{\ell}}},P_{1})=\sum_{p\mid{2\prod_{j=1}^{4}\alpha_{j}}}\sum_{\textbf{b}(p)\in\{0,1\}^{4}}\prod_{i=1}^{4}\mu(p^{b(p)_{i}})r_{X^{\prod_{p\mid 2\alpha_{j}}p^{c_{p}\cdot\textbf{b}(p)}\cdot\textbf{d}\cdot\boldsymbol{\ell}}}(h).$
We decompose
$r_{X^{\boldsymbol{d\ell}}}(h)=a_{E_{X^{\boldsymbol{d\ell}}}}(h)+(r_{X^{\boldsymbol{d\ell}}}(h)-a_{E_{X^{\boldsymbol{d\ell}}}}(h)).$ (5.5)
Then we have
$\begin{split}S_{\textbf{c},h}(A_{l},P_{\delta,\Delta_{\alpha}})\geq\sum_{{\textbf{d}\in S_{\prod_{p\mid 2\prod_{j=1}^{4}\alpha_{j}}p^{c_{p}}}^{4}\atop{p\mid d_{j}\implies p\in P_{\delta,\Delta_{\alpha}}}}}\Lambda_{d_{1}}^{-}\lambda_{d_{2}}^{+}\lambda_{d_{3}}^{+}\lambda_{d_{4}}^{+}(a_{E_{X^{\boldsymbol{d\ell}}}}(h)\\ +(r_{X^{\boldsymbol{d\ell}}}(h)-a_{E_{X^{\boldsymbol{d\ell}}}}(h))).\end{split}$ (5.6)
Since $\mid\Lambda_{d}^{-}\mid\leq 7$ pulling the absolute value inside the second sum in (5.4) yields the last term on the right-hand side of the proposition, and we henceforth ignore it.
Writing
$C_{p}:=\sum_{\textbf{b}(p)\in\{0,1\}^{4}}\prod_{i=1}^{4}\mu(p^{b(p)_{i}})(\frac{1}{p^{b(p)_{1}+b(p)_{2}+b(p)_{3}+b(p)_{4}}}\frac{b_{p}(h,\lambda_{p^{(b(p)_{1},b(p)_{2},b(p)_{3},b(p)_{4})}},0)}{b_{p}(h,\lambda_{\textbf{1}},0)})^{c_{p}},$
we can conclude that
$\begin{split}\sum_{{\textbf{d}\in S_{\prod_{p\mid 2\prod_{j=1}^{4}\alpha_{j}}p^{c_{p}}}^{4}\atop{p\mid d_{j}\implies p\in P_{\delta,\Delta_{\alpha}}}}}\Lambda_{d_{1}}^{-}\lambda_{d_{2}}^{+}\lambda_{d_{3}}^{+}\lambda_{d_{4}}^{+}a_{E_{X^{\boldsymbol{d\ell}}}}(h)=\prod_{p\mid 2\prod\alpha_{j}}C_{p}\sum_{{\textbf{d}\in S_{1}^{4}}\atop{p\mid d_{j}\implies p\in P_{\delta,\Delta_{\alpha}}}}\Lambda_{d_{1}}^{-}\lambda_{d_{2}}^{+}\lambda_{d_{3}}^{+}\lambda_{d_{4}}^{+}a_{E_{X^{\boldsymbol{d\ell}}}}(h).\end{split}$ (5.7)
Thus we may bound the main term from below by
$\prod_{p\mid 2\prod\alpha_{j}}C_{p}M_{h}^{-}(z_{0})a_{E_{X^{\boldsymbol{d\ell}}}}(h).$ (5.8)
(2) Using the other inequality yields the claim by the same steps as in (1). ∎
In order to prove Proposition 5.1, we need to get bounds for the main term.
Proof of Proposition 5.1.
The error terms are of the same size in Proposition 5.2 (1) and (2), so it remains to bound the main terms. We first consider the main term from Proposition 5.2 (2). We separate the terms off which have $\gcd(d_{j},d_{k})>\Delta$ for some $j\neq k.$
$\sum_{d_{1}\mid P(z_{0})}\lambda_{d_{1}}^{+}\sum_{d_{2}\mid P(z_{0})}\lambda_{d_{2}}^{+}\sum_{d_{3}\mid P(z_{0})}\lambda_{d_{3}}^{+}\sum_{d_{4}\mid P(z_{0})\atop\gcd(d_{3},d_{4})>\Delta}\lambda_{d_{4}}^{+}\prod_{p\mid\textbf{d}}\beta_{X^{\textbf{d}},p}(h)$
$\leq\sum_{d_{1}\mid P(z_{0})}\sum_{d_{2}\mid P(z_{0})}\sum_{d_{3}\mid P(z_{0})}\sum_{d_{4}\mid P(z_{0})\atop\gcd(d_{3},d_{4})>\Delta}\prod_{p\mid\textbf{d}}\beta_{X^{\textbf{d}},p}(h)$
$\ll(\sum_{d_{1}\mid P(z_{0})}\frac{\tilde{w}(d_{1})}{d_{1}})^{2}\sum_{\epsilon\geq\Delta}(\sum_{d_{2}\mid P(z_{0})\atop\epsilon\mid d_{2}}\frac{\tilde{w}(d_{2})}{d_{2}})^{2}$
$\ll\prod_{p\in P_{z_{0},\Delta_{\alpha}}}(1+\frac{\tilde{w}(p)}{p})^{4}\sum_{\epsilon\geq\Delta}\mu(\epsilon)^{2}(\frac{\tilde{w}(\epsilon)}{\epsilon})^{2}$
By Lemma 3.7
$\sum_{\epsilon\geq\Delta}\mu(\epsilon)^{2}(\frac{\tilde{w}(\epsilon)}{\epsilon})^{2}$ $\leq\sum_{\epsilon\geq\Delta}(\frac{\epsilon}{\Delta})^{\frac{1}{2}}\mu(\epsilon)^{2}(\frac{\tilde{w}(\epsilon)}{\epsilon})^{2}\leq\Delta^{-\frac{1}{2}}\sum_{\epsilon\geq 1}\mu(\epsilon)^{2}(\frac{\tilde{w}(\epsilon)^{2}}{\epsilon^{\frac{3}{2}}})$
$=\Delta^{-\frac{1}{2}}\prod_{p}(1+\frac{\tilde{w}(p)^{2}}{p^{\frac{3}{2}}})\leq\Delta^{-\frac{1}{2}}\prod_{p\mid n}(1+\frac{\tilde{w}(p)^{2}}{p^{\frac{3}{2}}})\prod_{p}(1+\frac{4}{p^{\frac{3}{2}}})$
$\ll\Delta^{-\frac{1}{2}}\prod_{p\mid n}(1+\frac{\tilde{w}(p)^{2}}{p^{\frac{3}{2}}})\ll\Delta^{-\frac{1}{2}}H(n).$
Then we have
$\prod_{p\in P_{z_{0},\Delta_{\alpha}}}(1+\frac{\tilde{w}(p)}{p})\leq\prod_{p\leq z_{0}}(1+\frac{\tilde{w}(p)}{p})\ll H(n)\log(z_{0})^{2}.$
Hence we may write $M_{h}^{+}(z_{0})$ as
$\sum_{d_{1}\mid P(z_{0})}\lambda_{d_{1}}^{+}\sum_{d_{2}\mid P(z_{0})}\lambda_{d_{2}}^{+}\sum_{d_{3}\mid P(z_{0})}\lambda_{d_{3}}^{+}\sum_{d_{4}\mid P(z_{0})\atop\gcd(d_{3},d_{4})\leq\Delta}\lambda_{d_{4}}^{+}\prod_{p\mid\textbf{d}}\beta_{X^{\textbf{d}},p}(h)+O(\Delta^{-\frac{1}{2}}H(n)^{5}\log(z_{0})^{8})$
By grouping those d with $\gcd(d_{i},d_{j})=u_{i,j}.$ We have
$\sum_{d_{1}\mid P(z_{0})}\lambda_{d_{1}}^{+}$ $\sum_{d_{2}\mid P(z_{0})}\lambda_{d_{2}}^{+}\sum_{d_{3}\mid P(z_{0})}\lambda_{d_{3}}^{+}\sum_{d_{4}\mid P(z_{0})\atop\gcd(d_{3},d_{4})\leq\Delta}\lambda_{d_{4}}^{+}\prod_{p\mid\textbf{d}}\beta_{X^{\textbf{d}},p}(h)$
$=\sum_{u_{1,2}\leq\Delta}\cdot\cdot\cdot\sum_{u_{3,4}\leq\Delta}\sum_{d_{1}\mid P(z_{0})}\lambda_{d_{1}}^{+}\sum_{d_{2}\mid P(z_{0})}\lambda_{d_{2}}^{+}\sum_{d_{3}\mid P(z_{0})}\lambda_{d_{3}}^{+}\sum_{d_{4}\mid P(z_{0})\atop\gcd(d_{3},d_{4})=u_{i,j}}\lambda_{d_{4}}^{+}\prod_{p\mid\textbf{d}}\beta_{X^{\textbf{d}},p}(h).$
Since $g(\textbf{d})$ only depends on $u_{i,j}$ by Lemma 3.8, we may write
$\sum_{u_{1,2}\leq\Delta}\cdot\cdot\cdot\sum_{u_{3,4}\leq\Delta}\sum_{d_{1}\mid P(z_{0})}\lambda_{d_{1}}^{+}\sum_{d_{2}\mid P(z_{0})}\lambda_{d_{2}}^{+}\sum_{d_{3}\mid P(z_{0})}\lambda_{d_{3}}^{+}\sum_{d_{4}\mid P(z_{0})\atop\gcd(d_{3},d_{4})=u_{i,j}}\lambda_{d_{4}}^{+}\prod_{p\mid\textbf{d}}\beta_{X^{\textbf{d}},p}(h)\\ =\sum_{u_{1,2}\leq\Delta}\cdot\cdot\cdot\sum_{u_{3,4}\leq\Delta}g(\textbf{d})(\sum_{d_{1}\mid P(z_{0})}\lambda_{d_{1}}^{+}\prod_{p\mid\textbf{d}_{1}}\beta_{X^{(p,1,1,1)},p}(h)\sum_{d_{2}\mid P(z_{0})}\lambda_{d_{2}}^{+}\prod_{p\mid\textbf{d}_{2}}\beta_{X^{(1,p,1,1)},p}(h)\\ \cdot\sum_{d_{3}\mid P(z_{0})}\lambda_{d_{3}}^{+}\prod_{p\mid\textbf{d}_{3}}\beta_{X^{(1,1,p,1)},p}(h)\sum_{d_{4}\mid P(z_{0})\atop\gcd(d_{3},d_{4})=u_{i,j}}\lambda_{d_{4}}^{+}\prod_{p\mid\textbf{d}_{4}}\beta_{X^{(1,1,1,p)},p}(h)).$
We rewrite the $\gcd$ condition on the inner sums by excluding divisibility by other primes using $\mu$ (or inclusion/exclusion). Letting $l_{i,j}$ run through all possible divisors of $P(z_{0})$ relatively prime to $u_{i,j}$, this gives
$\sum_{d_{1}\mid P(z_{0})}\lambda_{d_{1}}^{+}\prod_{p\mid\textbf{d}_{1}}\beta_{X^{(p,1,1,1)},p}(h)\sum_{d_{2}\mid P(z_{0})}\lambda_{d_{2}}^{+}\prod_{p\mid\textbf{d}_{2}}\beta_{X^{(1,p,1,1)},p}(h)\\ \cdot\sum_{d_{3}\mid P(z_{0})}\lambda_{d_{3}}^{+}\prod_{p\mid\textbf{d}_{3}}\beta_{X^{(1,1,p,1)},p}(h)\sum_{d_{4}\mid P(z_{0})\atop\gcd(d_{3},d_{4})=u_{i,j}}\lambda_{d_{4}}^{+}\prod_{p\mid\textbf{d}_{4}}\beta_{X^{(1,1,1,p)},p}(h)$
$\sum_{l_{i,j}\atop 1\leq i,j\leq 4}\mu(l_{i,j})\sum_{d_{1}\mid P(z_{0})\atop lcm(u_{1,2}l_{1,2},u_{1,3}l_{1,3},u_{1,4}l_{1,4})}\lambda_{d_{1}}^{+}\prod_{p\mid\textbf{d}_{1}}\beta_{X^{(p,1,1,1)},p}(h)\\ \cdot\cdot\cdot\sum_{d_{4}\mid P(z_{0})\atop lcm(u_{1,4}l_{1,4},u_{2,4}l_{2,4},u_{3,4}l_{3,4})}\lambda_{d_{4}}^{+}\prod_{p\mid\textbf{d}_{4}}\beta_{X^{(1,1,1,p)},p}(h).$
By Lemma 3.7 we have
$\sum_{d_{1}\mid P(z_{0})\atop\delta\mid d_{1}}\lambda_{d_{1}}^{+}\prod_{p\mid\textbf{d}_{1}}\beta_{X^{(p,1,1,1)},p}(h)\leq\prod_{p\mid\delta}\beta_{X^{(\delta,1,1,1)},p}(h)\cdot\prod_{p\mid P(z_{0})}(1+\frac{2}{p})\\ \ll\frac{2^{w(\delta)}}{\varphi(\delta)}\log(z_{0})^{2}\ll\frac{\sigma(\delta)^{2}}{\delta}\log(z_{0})^{2}.$
Then we want to bound
$\sum_{l_{1,2},\cdot\cdot\cdot,l_{3,4}\leq D_{0}\atop\Delta^{B}<l_{1,2}}\frac{\mu(l_{1,2})^{2}\cdot\cdot\cdot\mu(l_{3,4})^{2}\sigma_{0}(l_{1,2})^{4}\cdot\cdot\cdot\sigma_{0}(l_{3,4})^{4}\gcd(l_{1,2},l_{1,3})\cdot\cdot\cdot\gcd(l_{2,4},l_{3,4})}{l_{1,2}^{2}\cdot\cdot\cdot l_{3,4}^{2}\gcd(l_{1,2},l_{1,3},l_{1,4})\cdot\cdot\cdot\gcd(l_{1,4},l_{2,4},l_{3,4})}$
We then bound
$\sum_{l_{3,4}\leq D_{0}}\frac{\mu(l_{3,4})^{2}\sigma_{0}(l_{3,4})^{4}\gcd(l_{1,3},l_{3,4})\gcd(l_{1,4},l_{3,4})\gcd(l_{2,4},l_{3,4})}{l_{3,4}^{2}\gcd(l_{1,3},l_{2,3},l_{3,4})\gcd(l_{1,4},l_{2,4},l_{3,4})}\\ \leq\prod_{p\mid l_{1,3}l_{2,3}l_{1,4}l_{2,4}}(2^{4}+1)\prod_{p}(1+\frac{2^{4}}{p^{2}})\ll(\sigma_{0}(l_{1,3})\sigma_{0}(l_{1,4})\sigma_{0}(l_{2,3})\sigma_{0}(l_{2,4}))^{5}$
Then bound
$\sum_{l_{2,4}\leq D_{0}}\frac{\mu(l_{2,4})^{2}\sigma_{0}(l_{2,4})^{9}\gcd(l_{1,2},l_{2,4})\gcd(l_{2,3},l_{2,4})\gcd(l_{1,4},l_{2,4})}{l_{2,4}^{2}\gcd(l_{1,2},l_{2,3},l_{2,4})}\\ \leq\prod_{p\mid l_{1,3}l_{2,3}l_{1,4}}(2^{9}+1)\prod_{p}(1+\frac{2^{9}}{p^{2}})\ll(\sigma_{0}(l_{1,2})\sigma_{0}(l_{2,3})\sigma_{0}(l_{1,4}))^{10}$
Then bound
$\sum_{l_{1,4}\leq D_{0}}\frac{\mu(l_{1,4})^{2}\sigma_{0}(l_{1,4})^{19}\gcd(l_{1,2},l_{1,4})\gcd(l_{1,3},l_{1,4})}{l_{1,4}^{2}\gcd(l_{1,2},l_{1,3},l_{1,4})}\\ \leq\prod_{p\mid l_{1,2}l_{1,3}}(2^{19}+1)\prod_{p}(1+\frac{2^{19}}{p^{2}})\ll(\sigma_{0}(l_{1,2})\sigma_{0}(l_{1,3}))^{20}$
Then bound
$\sum_{l_{2,3}\leq D_{0}}\frac{\mu(l_{2,3})^{2}\sigma_{0}(l_{2,3})^{19}\gcd(l_{1,2},l_{2,3})\gcd(l_{1,3},l_{2,3})}{l_{2,3}^{2}}\\ \leq\prod_{p\mid l_{1,2}}(2^{19}+1)\prod_{p}(1+\frac{2^{19}}{p^{2}})\ll\sigma_{0}(l_{1,2})^{20}$
Finally the sum over $i_{1,2}$ is bounded by
$\sum_{l_{1,2}>\Delta^{B}}\frac{\mu(l_{1,2})^{2}\sigma_{0}(l_{1,2})^{54}}{l_{1,2}^{2}}\ll_{\epsilon}\Delta^{\epsilon-B}.$
Therefore the contribution from $l_{i,j}$ with at least one of them $\Delta^{B}$ is
$\ll_{\epsilon}\Delta^{\epsilon-B}\log(z_{0})^{8}.$
And we finally use [2, Lemma 11] to finish the proof for the upper bound. Since we have taken the absolute value at each step, the proof for the lower bound follows by the same argument, using Proposition 5.2 (1) as a starting point. ∎
6 Sieving and proof of theorem 1.1
We employ a second sieve to prove Theorem 1.1.
Proof of Theorem 1.1.
Suppose that $z_{0}<z$ and set $P_{h}(z_{0},z):=\prod_{z_{0}<p\leq z\atop p\mid\prod_{j=1}^{4}\alpha_{j}}p.$ Then we have
$S_{c,h}(A_{\textbf{1},z})$ $=\sum_{F_{\textbf{c,h}}(A_{\textbf{1}},z_{0})}\prod_{j=1}^{4}\sum_{l_{j}\mid\gcd(x_{j},P_{h}(z_{0},z))}\mu(l_{j})$
$=\sum_{F_{\textbf{c,h}}(A_{\textbf{1}},z_{0})}\sum_{l_{1}\mid P_{h}(z_{0},z)\atop l_{1}\mid x_{1}}\mu(l_{1})\sum_{l_{2}\mid P_{h}(z_{0},z)\atop l_{2}\mid x_{2}}\mu(l_{2})\sum_{l_{3}\mid P_{h}(z_{0},z)\atop l_{3}\mid x_{3}}\mu(l_{3})\sum_{l_{4}\mid P_{h}(z_{0},z)\atop l_{4}\mid x_{4}}\mu(l_{4})$
$\geq\sum_{F_{\textbf{c,h}}(A_{\textbf{1}},z_{0})}\sum_{l_{1}\mid P_{h}(z_{0},z)\atop l_{1}\mid x_{1}}\Lambda_{l_{1}}^{-}\sum_{l_{2}\mid P_{h}(z_{0},z)\atop l_{2}\mid x_{2}}\lambda_{l_{2}}^{+}\sum_{l_{3}\mid P_{h}(z_{0},z)\atop l_{3}\mid x_{3}}\lambda_{l_{3}}^{+}\sum_{l_{4}\mid P_{h}(z_{0},z)\atop l_{4}\mid x_{4}}\lambda_{l_{4}}^{+}$
$=\sum_{l_{1}\mid P_{h}(z_{0},z)}\Lambda_{l_{1}}^{-}\sum_{l_{2}\mid P_{h}(z_{0},z)}\lambda_{l_{2}}^{+}\sum_{l_{3}\mid P_{h}(z_{0},z)}\lambda_{l_{3}}^{+}\sum_{l_{4}\mid P_{h}(z_{0},z)}\lambda_{l_{4}}^{+}S_{\textbf{c},h}(A_{\boldsymbol{\ell}},z_{0}).$
We then plug in Proposition 5.1 in the inner sum to obtain
$S_{\textbf{c},h}(A_{\textbf{1}},z)\geq\sum_{\ell_{1}\mid P_{h}(z_{0},z)}\Lambda_{\ell_{1}}^{-}\sum_{\ell_{2}\mid P_{h}(z_{0},z)}\lambda_{\ell_{2}}^{+}\sum_{\ell_{3}\mid P_{h}(z_{0},z)}\lambda_{\ell_{3}}^{+}\sum_{\ell_{4}\mid P_{h}(z_{0},z)}\lambda_{\ell_{4}}^{+}(W_{c,h}(n^{\delta})+\\ O_{\epsilon}(\Delta^{-\frac{1}{2}}H(n)^{5}\log(n^{\delta})+\Delta^{\epsilon-B}\log(n^{\delta})^{8}+\Delta^{C}e^{-s_{0}}))a_{E_{X^{\boldsymbol{d\ell}}}}(h)\\ +O(\sum_{{\boldsymbol{\ell}\in S_{\prod_{p\mid 2\prod_{j=1}^{4}\alpha_{j}}p^{c_{p}}}^{4}\atop{p\mid\ell_{j}\implies p\mid P_{h}(z_{0},z)\atop\ell_{j}\leq D}}}\sum_{{\textbf{d}\in S_{\prod_{p\mid 2\prod_{j=1}^{4}\alpha_{j}}p^{c_{p}}}^{4}\atop{p\mid d_{j}\implies p\in P_{\delta,\Delta_{\alpha}}}}\atop d_{j}\leq D_{0}}\mid r_{X^{{\textbf{d}}\cdot{\boldsymbol{\ell}}}}(h)-a_{E_{X^{\boldsymbol{d\ell}}}}(h)\mid)).$
We first bound the last O-term. Lemma 4.2 implies that
$\sum_{{\boldsymbol{\ell}\in S_{\prod_{p\mid 2\prod_{j=1}^{4}\alpha_{j}}p^{c_{p}}}^{4}\atop{p\mid l_{j}\implies p\mid P_{h}(z_{0},z)\atop l_{j}\leq D}}}$ $\sum_{{\textbf{d}\in S_{\prod_{p\mid 2\prod_{j=1}^{4}\alpha_{j}}p^{c_{p}}}^{4}\atop{p\mid d_{j}\implies p\in P_{\delta,\Delta_{\alpha}}}}\atop d_{j}\leq D_{0}}\mid r_{X^{{\textbf{d}}\cdot{\boldsymbol{\ell}}}}(h)-a_{E_{X^{\boldsymbol{d\ell}}}}(h)\mid$
$\ll\sum_{{\boldsymbol{\ell}\in S_{\prod_{p\mid 2\prod_{j=1}^{4}\alpha_{j}}p^{c_{p}}}^{4}\atop{p\mid l_{j}\implies p\mid P_{h}(z_{0},z)\atop l_{j}\leq D}}}\sum_{{\textbf{d}\in S_{\prod_{p\mid 2\prod_{j=1}^{4}\alpha_{j}}p^{c_{p}}}^{4}\atop{p\mid d_{j}\implies p\in P_{\delta,\Delta_{\alpha}}}}\atop d_{j}\leq D_{0}}N_{\alpha,\boldsymbol{d\ell}}^{\frac{5}{2}+\epsilon}h^{\frac{1}{2}+\epsilon}$
where $N_{\alpha,\boldsymbol{d\ell}}$ is the level of $f_{\boldsymbol{\alpha}}(\boldsymbol{d\ell}\textbf{x}).$ Since $\ell_{j}\leq D$, $d_{j}\leq D_{0},$ $N_{\alpha,\boldsymbol{d\ell}}\ll_{m}\prod_{j=1}^{4}d_{j}^{2}l_{j}^{2}\alpha_{j},$ we have
$\sum_{{\boldsymbol{\ell}\in S_{\prod_{p\mid 2\prod_{j=1}^{4}\alpha_{j}}p^{c_{p}}}^{4}\atop{p\mid\ell_{j}\implies p\mid P_{h}(z_{0},z)\atop\ell_{j}\leq D}}}$ $\sum_{{\textbf{d}\in S_{\prod_{p\mid 2\prod_{j=1}^{4}\alpha_{j}}p^{c_{p}}}^{4}\atop{p\mid d_{j}\implies p\in P_{\delta,\Delta_{\alpha}}}}\atop d_{j}\leq D_{0}}N_{\alpha,\boldsymbol{d\ell}}^{\frac{11}{4}+\epsilon}h^{\frac{3}{5}}\ll_{m}(DD_{0})^{4}(DD_{0})^{22+\epsilon}(\prod_{j=1}^{4}\alpha_{j})^{\frac{11}{4}+\epsilon}h^{\frac{1}{2}+\epsilon}$ (6.1)
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad=(DD_{0})^{26+\epsilon}(\prod_{j=1}^{4}\alpha_{j})^{\frac{11}{4}+\epsilon}h^{\frac{1}{2}+\epsilon}.$
[…]
$(\Delta^{-\frac{1}{2}}H(n)^{5}\log(n^{\delta})+\Delta^{\epsilon-B}\log(n^{\delta})^{8}+\Delta^{C}e^{-s_{0}})\sum_{{\boldsymbol{\ell}\in S_{\prod_{p\mid 2\prod_{j=1}^{4}\alpha_{j}}p^{c_{p}}}^{4}\atop{p\mid\ell_{j}\implies p\mid P_{h}(z_{0},z)\atop l_{j}\leq D}}}a_{E_{X^{\boldsymbol{d\ell}}}}(h)$ (6.2)
$\ll(\Delta^{-\frac{1}{2}}H(n)^{5}\log(n^{\delta})+\Delta^{\epsilon-B}\log(n^{\delta})^{8}+\Delta^{C}e^{-s_{0}})(\frac{\log(z)}{\log(z_{0})})^{6}a_{E_{X^{\boldsymbol{d\ell}}}}(h).$
For the main term, we write
$\sum_{\ell_{1}\mid P_{h}(z_{0},z)}\Lambda_{\ell_{1}}^{-}\sum_{\ell_{2}\mid P_{h}(z_{0},z)}\lambda_{\ell_{2}}^{+}\sum_{\ell_{3}\mid P_{h}(z_{0},z)}\lambda_{\ell_{3}}^{+}\sum_{\ell_{4}\mid P_{h}(z_{0},z)}\lambda_{\ell_{4}}^{+}\prod_{p\mid\boldsymbol{\ell}}\beta_{X^{\boldsymbol{\ell}},p}(h)$ (6.3)
$\sum_{\ell_{1}\mid P_{h}(z_{0},z)}\Lambda_{\ell_{1}}^{-}\sum_{\ell_{2}\mid P_{h}(z_{0},z)}\lambda_{\ell_{2}}^{+}\sum_{\ell_{3}\mid P_{h}(z_{0},z)}\lambda_{\ell_{3}}^{+}\sum_{\ell_{4}\mid P_{h}(z_{0},z)\atop\gcd(\ell_{i},\ell_{j})=1}\lambda_{\ell_{4}}^{+}\prod_{p\mid\boldsymbol{\ell}}\beta_{X^{\boldsymbol{\ell}},p}(h)\\ +\sum_{\boldsymbol{\ell}\mid P_{h}(z_{0},z)^{4}\atop\exists(j,k):\gcd(\ell_{j},\ell_{k})\neq 1}\Lambda_{\ell_{1}}^{-}\lambda_{\ell_{2}}^{+}\lambda_{\ell_{3}}^{+}\lambda_{\ell_{4}}^{+}\prod_{p\mid\boldsymbol{\ell}}\beta_{X^{\boldsymbol{\ell}},p}(h)$
$\sum_{\ell_{1}\mid P_{h}(z_{0},z)}\Lambda_{\ell_{1}}^{-}\prod_{p\mid\ell_{1}}\beta_{X^{(\ell_{1},1,1,1)},p}(h)\sum_{\ell_{2}\mid P_{h}(z_{0},z)}\lambda_{\ell_{2}}^{+}\prod_{p\mid\ell_{2}}\beta_{X^{(1,\ell_{1},1,1)},p}(h)\\ \cdot\sum_{\ell_{3}\mid P_{h}(z_{0},z)}\lambda_{\ell_{3}}^{+}\prod_{p\mid\ell_{3}}\beta_{X^{(1,1,\ell_{1},1)},p}(h)\sum_{\ell_{4}\mid P_{h}(z_{0},z)}\lambda_{\ell_{4}}^{+}\prod_{p\mid\ell_{4}}\beta_{X^{(1,1,1,\ell_{1})},p}(h)\\ +\sum_{\boldsymbol{\ell}\mid p_{h}(z_{0},z)^{4}\atop\exists(j,m):\gcd(\ell_{j},\ell_{k})\neq 1}\Lambda_{\ell_{1}}^{-}\lambda_{\ell_{2}}^{+}\lambda_{\ell_{3}}^{+}\lambda_{\ell_{4}}^{+}(\prod_{p\mid\boldsymbol{\ell}}\beta_{X^{\boldsymbol{\ell}},p}(h)-\prod_{p\mid\ell_{1}}\beta_{X^{(\ell_{1},1,1,1)},p}(h)\prod_{p\mid\ell_{2}}\beta_{X^{(1,\ell_{1},1,1)},p}(h)\\ \cdot\prod_{p\mid\ell_{3}}\beta_{X^{(1,1,\ell_{1},1)},p}(h)\prod_{p\mid\ell_{4}}\beta_{X^{(1,1,1,\ell_{1})},p}(h)).$
Abbreviating $\beta_{\ell_{1},g}:=\prod_{p\mid\gcd(\ell_{1},g)}\beta_{X^{(\ell_{1},1,1,1)},p}(h),...,\beta_{\ell_{4},g}:=\prod_{p\mid\gcd(\ell_{4},g)}\beta_{X^{(1,1,1,\ell_{1})},p}(h)$ Then we can bound the last sum as
$\ll\sum_{g\mid P_{h}(z_{0},z)\atop g\neq 1}\sum_{\begin{subarray}{c}\boldsymbol{\ell}\mid P_{h}(z_{0},z)^{4}\\ g_{1}=\gcd(\ell_{1},\ell_{2})\mid g\\ ...\\ g_{6}=\gcd(\ell_{3},\ell_{4})\mid g)\end{subarray}}\prod_{p\mid\boldsymbol{\ell}\atop p\nmid g}\beta_{X^{\boldsymbol{\ell}},p}(h)\mid\prod_{p\mid g}\beta_{X^{\boldsymbol{\ell}},p}(h)-\beta_{\ell_{1},g}\beta_{\ell_{2},g}\beta_{\ell_{3},g}\beta_{\ell_{4},g}\mid.$ (6.4)
We then bound the absolute value by
$\mid\prod_{p\mid g}\beta_{X^{\boldsymbol{\ell}},p}(h)-\beta_{\ell_{1},g}\beta_{\ell_{2},g}\beta_{\ell_{3},g}\beta_{\ell_{4},g}\mid\leq\max\{\prod_{p\mid g}\beta_{X^{\boldsymbol{\ell}},p}(h),\beta_{\ell_{1},g}\beta_{\ell_{2},g}\beta_{\ell_{3},g}\beta_{\ell_{4},g}\}$ (6.5)
$\leq\prod_{p\mid g}\max\{\max\{\beta_{X^{\boldsymbol{\ell}},p}(h)\},\max_{i}\{\beta_{\ell_{i},g}\}^{\#\{j:p\mid\ell_{j}\}}\}.$
The product after sum becomes
$\leq\prod_{p\mid l\atop\#\{j:p\mid\ell_{j}\}=1}\max_{i}\{\beta_{\ell_{i},p}(h)\}\cdot\prod_{p\mid g\atop\#\{j:p\mid\ell_{j}\}=2}\max\{\max\{\beta_{X^{(p,p,1,1)},p}(h),...,\beta_{X^{(1,1,p,p)},p}(h)\},\max_{i}\{\beta_{\ell_{i},g}\}^{2}\}$ (6.6)
$\cdot\prod_{p\mid g\atop\#\{j:p\mid\ell_{j}\}=3}\max\{\max\{\beta_{X^{(p,p,p,1)},p}(h),...,\beta_{X^{(1,p,p,p)},p}(h)\},\max_{i}\{\beta_{\ell_{i},g}\}^{3}\}$
$\cdot\prod_{p\mid g\atop\#\{j:p\mid\ell_{j}\}=4}\max\{\beta_{X^{(p,p,p,p)},p}(h)\},\max_{i}\{\beta_{\ell_{i},g}\}^{4}\}.$
Without loss of generality, taking $z_{0}$ sufficiently large, we may assume that $p\nmid 2(m-2)(m-4)$ for every $p\mid\boldsymbol{\ell}.$ Thus, using Lemma 3.9, we may bound (6.6) against
$\leq\prod_{p\mid{\boldsymbol{\ell}}\atop\#\{j:p\mid\ell_{j}\}=1}\frac{4}{p}\cdot\prod_{p\mid g\atop\#\{j:p\mid\ell_{j}\}=2}\frac{16}{p^{2}}\cdot\prod_{p\mid g\atop\#\{j:p\mid\ell_{j}\}=3}\frac{64}{p^{3}}\cdot\prod_{\begin{subarray}{c}p\mid{\boldsymbol{\ell}}\\ \#\{j:p\mid\ell_{j}\}=1\\ p\nmid n\end{subarray}}\frac{256}{p^{4}}\cdot\prod_{\begin{subarray}{c}p\mid{\boldsymbol{\ell}}\\ \#\{j:p\mid\ell_{j}\}=1\\ p\mid n\end{subarray}}\frac{1}{(p-1)^{2}(p+1)}$
$=\frac{1}{l_{1}l_{2}l_{3}l_{4}}\prod_{p\mid{\boldsymbol{\ell}}\atop\#\{j:p\mid\ell_{j}\}=1}4\prod_{p\mid g\atop\#\{j:p\mid\ell_{j}\}=2}4\prod_{p\mid g\atop\#\{j:p\mid\ell_{j}\}=3}16\prod_{\begin{subarray}{c}p\mid{\boldsymbol{\ell}}\\ \#\{j:p\mid\ell_{j}\}=1\\ p\nmid n\end{subarray}}64\prod_{\begin{subarray}{c}p\mid{\boldsymbol{\ell}}\\ \#\{j:p\mid\ell_{j}\}=1\\ p\mid n\end{subarray}}\frac{p^{4}}{4(p-1)^{2}(p+1)}$
$\leq\frac{4^{w^{(\ell_{1}\ell_{2}\ell_{3}\ell_{4})}}64^{w(g)}}{\ell_{1}\ell_{2}\ell_{3}\ell_{4}}\gcd(g,n,\ell_{1},\ell_{2},\ell_{3},\ell_{4})$
$\leq\frac{4^{w(\ell_{1})}}{l_{1}}\frac{4^{w(\ell_{2})}}{l_{2}}\frac{4^{w(\ell_{3})}}{l_{3}}\frac{4^{w(\ell_{4})}}{l_{4}}64^{w(g)}\gcd(g,n,\ell_{1},\ell_{2},\ell_{3},\ell_{4}).$
Hence we can bound (6.4) by
$\sum_{g\mid P_{h}(z_{0},z)}\sum_{\begin{subarray}{c}\boldsymbol{\ell}\mid P_{h}(z_{0},z)^{4}\\ g_{1,2}=\gcd(\ell_{1},\ell_{2})\mid g\\ ...\\ g_{3,4}=\gcd(\ell_{3},\ell_{4})\mid g\\ lcm(g_{1,2},...,g_{3,4})=g\end{subarray}}\frac{4^{w(\ell_{1})}}{\ell_{1}}\frac{4^{w(\ell_{2})}}{\ell_{2}}\frac{4^{w(\ell_{3})}}{\ell_{3}}\frac{4^{w(\ell_{4})}}{\ell_{4}}64^{w(g)}\gcd(g,n,\ell_{1},\ell_{2},\ell_{3},\ell_{4}).$ (6.7)
Let $g_{1}:=\text{lcm}(g_{1,2},g_{1,3},g_{1,4}),...,g_{4}:=\text{lcm}(g_{1,4},g_{2,4},g_{3,4}),$ and since $g_{j}\mid\ell_{j},$ by a change of variable $\ell_{j}\xrightarrow{}g_{j}\ell_{j},$ (6.7) become
$\sum_{\begin{subarray}{c}g\mid P_{h}(z_{0},z)\\ g\neq 1\\ g_{1},g_{2},g_{3},g_{4}\mid g\\ \text{lcm}(\gcd(g_{1},g_{2}),...,\gcd(g_{3},g_{4}))=g\end{subarray}}\frac{4^{w(g_{1})}}{g_{1}}\frac{4^{w(g_{2})}}{g_{2}}\frac{4^{w(g_{3})}}{g_{3}}\frac{4^{w(g_{4})}}{g_{4}}64^{w(g)}\gcd(n,g_{1},g_{2},g_{3},g_{4}).\\ \cdot\sum_{\begin{subarray}{c}\boldsymbol{\ell}\mid P_{h}(z_{0},z)^{4}\\ g_{1,2}=\gcd(\ell_{1},\ell_{2})\mid g\\ ...\\ g_{3,4}=\gcd(\ell_{3},\ell_{4})\mid g\\ lcm(g_{1,2},...,g_{3,4})=g\end{subarray}}\frac{4^{w(\ell_{1})}}{\ell_{1}}\frac{4^{w(\ell_{2})}}{\ell_{2}}\frac{4^{w(\ell_{3})}}{\ell_{3}}\frac{4^{w(\ell_{4})}}{\ell_{4}}\\ \leq\sum_{\begin{subarray}{c}g\mid P_{h}(z_{0},z)\\ g\neq 1\\ g_{1},g_{2},g_{3},g_{4}\mid g\\ \text{lcm}(\gcd(g_{1},g_{2}),...,\gcd(g_{3},g_{4}))=g\end{subarray}}\prod_{p\mid g}\frac{4^{4}64p^{\text{min}\{\text{ord}_{p}(g_{1}),\text{ord}_{p}(g_{2}),\text{ord}_{p}(g_{3}),\text{ord}_{p}(g_{4})\}}}{p^{\text{ord}_{p}(g_{1})+\text{ord}_{p}(g_{2})+\text{ord}_{p}(g_{3})+\text{ord}_{p}(g_{4})}}\\ \cdot\sum_{\boldsymbol{\ell}\mid P_{h}(z_{0},z)^{4}}\frac{4^{w(\ell_{1})}}{\ell_{1}}\frac{4^{w(\ell_{2})}}{\ell_{2}}\frac{4^{w(\ell_{3})}}{\ell_{3}}\frac{4^{w(\ell_{4})}}{\ell_{4}}$
We then note that for $p\mid g$
$2\leq\text{ord}_{p}(g_{1})+\text{ord}_{p}(g_{2})+\text{ord}_{p}(g_{3})+\text{ord}_{p}(g_{4})\\ -\text{min}\{\text{ord}_{p}(g_{1}),\text{ord}_{p}(g_{2}),\text{ord}_{p}(g_{3}),\text{ord}_{p}(g_{4})\}\leq 3$
because $\text{lcm}(g_{1,2},...,g_{3,4})=g.$ Writing
$\sum_{\boldsymbol{\ell}\mid P_{h}(z_{0},z)^{4}}\frac{4^{w(\ell_{1})}}{\ell_{1}}\frac{4^{w(\ell_{2})}}{\ell_{2}}\frac{4^{w(\ell_{3})}}{\ell_{3}}\frac{4^{w(\ell_{4})}}{\ell_{4}}=\prod_{p\mid P_{h}(z_{0},z)}(1+\frac{4}{p})^{4}\\ \leq\prod_{p\mid P_{h}(z_{0},z)}(1+\frac{1}{p})^{16}\ll(\frac{\log(z)}{\log(z_{0})})^{16}.$
And writing
$\sum_{\begin{subarray}{c}g\mid P_{h}(z_{0},z)\\ g\neq 1\\ g_{1},g_{2},g_{3},g_{4}\mid g\\ \text{lcm}(\gcd(g_{1},g_{2}),...,\gcd(g_{3},g_{4}))=g\end{subarray}}\prod_{p\mid g}\frac{4^{3}64p^{\text{min}\{\text{ord}_{p}(g_{1}),\text{ord}_{p}(g_{2}),\text{ord}_{p}(g_{3}),\text{ord}_{p}(g_{4})\}}}{p^{\text{ord}_{p}(g_{1})+\text{ord}_{p}(g_{2})+\text{ord}_{p}(g_{3})+\text{ord}_{p}(g_{4})}}\\ \leq\sum_{\begin{subarray}{c}g\mid P_{h}(z_{0},z)\\ g\neq 1\\ g_{1},g_{2},g_{3},g_{4}\mid g\end{subarray}}\prod_{p\mid g}\frac{4^{3}64}{p^{2}}=\sum_{\begin{subarray}{c}g\mid P_{h}(z_{0},z)\\ g\neq 1\end{subarray}}\sigma_{0}(g)^{4}\prod_{p\mid g}\frac{4^{3}64}{p^{2}}\ll_{\epsilon}\sum_{g\geq z_{0}}g^{\epsilon-2}\ll_{\epsilon}z_{0}^{\epsilon-1}$
Thus we have the following bound
$\sum_{\boldsymbol{\ell}\mid p_{h}(z_{0},z)^{4}\atop\exists(j,m):\gcd(\ell_{j},\ell_{k})\neq 1}\Lambda_{\ell_{1}}^{-}\lambda_{\ell_{2}}^{+}\lambda_{\ell_{3}}^{+}\lambda_{\ell_{4}}^{+}(\prod_{p\mid\boldsymbol{\ell}}\beta_{X^{\boldsymbol{\ell}},p}(h)-\prod_{p\mid\ell_{1}}\beta_{X^{(\ell_{1},1,1,1)},p}(h)\prod_{p\mid\ell_{2}}\beta_{X^{(1,\ell_{1},1,1)},p}(h)\\ \cdot\prod_{p\mid\ell_{3}}\beta_{X^{(1,1,\ell_{1},1)},p}(h)\prod_{p\mid\ell_{4}}\beta_{X^{(1,1,1,\ell_{1})},p}(h)).\\ \ll_{\epsilon}z_{0}^{\epsilon-1}(\frac{\log(z)}{\log(z_{0})})^{16}$
Plugging in back to (6.3), we have
$\sum_{\ell_{1}\mid P_{h}(z_{0},z)}\Lambda_{\ell_{1}}^{-}\sum_{\ell_{2}\mid P_{h}(z_{0},z)}\lambda_{\ell_{2}}^{+}\sum_{\ell_{3}\mid P_{h}(z_{0},z)}\lambda_{\ell_{3}}^{+}\sum_{\ell_{4}\mid P_{h}(z_{0},z)}\lambda_{\ell_{4}}^{+}\prod_{p\mid\boldsymbol{\ell}}\beta_{X^{\boldsymbol{\ell}},p}(h)$ (6.8)
$\sum_{\ell_{1}\mid P_{h}(z_{0},z)}\Lambda_{\ell_{1}}^{-}\prod_{p\mid\ell_{1}}\beta_{X^{(\ell_{1},1,1,1)},p}(h)\sum_{\ell_{2}\mid P_{h}(z_{0},z)}\lambda_{\ell_{2}}^{+}\prod_{p\mid\ell_{2}}\beta_{X^{(1,\ell_{1},1,1)},p}(h)\\ \cdot\sum_{\ell_{3}\mid P_{h}(z_{0},z)}\lambda_{\ell_{3}}^{+}\prod_{p\mid\ell_{3}}\beta_{X^{(1,1,\ell_{1},1)},p}(h)\sum_{\ell_{4}\mid P_{h}(z_{0},z)}\lambda_{\ell_{4}}^{+}\prod_{p\mid\ell_{4}}\beta_{X^{(1,1,1,\ell_{1})},p}(h)\\ +O_{\epsilon}(z_{0}^{\epsilon-1}log(z)^{16})$
By [3, (6.32)], we have
$\sum_{\ell_{2}\mid P_{h}(z_{0},z)}\lambda_{\ell_{2}}^{+}\prod_{p\mid\ell_{2}}\beta_{X^{(1,\ell_{1},1,1)},p}(h)\geq\prod_{z_{0}<p\leq z\atop p\mid\prod_{j=1}^{4}\alpha_{j}}(1-\beta_{X^{(1,p,1,1)},p}(h))$ (6.9)
We then bound (for $z_{0}\geq 5$)
$\prod_{z_{0}<p\leq z\atop p\mid\prod_{j=1}^{4}\alpha_{j}}(1-\beta_{X^{(1,p,1,1)},p}(h))\geq\prod_{z_{0}<p\leq z}(1-\frac{4}{p})\geq\frac{1}{\zeta(4)}\prod_{z_{0}<p\leq z}(1-\frac{1}{p})^{4}.$
Using [6, (3.30) and (3.26)], for $z_{0}$ sufficiently large we obtain from (6.8) that
$\sum_{\ell_{2}\mid P_{h}(z_{0},z)}\lambda_{\ell_{2}}^{+}\prod_{p\mid\ell_{2}}\beta_{X^{(1,\ell_{1},1,1)},p}(h)\gg(\frac{\log(z_{0})}{\log(z)})^{4}$ (6.10)
Taking $\kappa=4$ and
$K:=\zeta(4)(1+\frac{1}{2\log(z)^{2}})^{4}(1+\frac{1}{\log(z_{0})^{2}})^{4}$
We can conclude that for sufficiently large $z_{0}$
$\prod_{z_{0}<p\leq z\atop p\mid\prod_{j=1}^{4}\alpha_{j}}(1-\beta_{X^{(1,p,1,1)},p}(h))\geq K^{-1}(\frac{\log(z_{0})}{\log(z)})^{4}$ (6.11)
By choosing $z_{0}$ sufficiently large, we can take $K$ arbitrarily close to $\zeta(4)$ from above, say $K<1.645.$ By [3, Theorem 6.1], we hence obtain
$\sum_{\ell_{1}\mid P_{h}(z_{0},z)}\Lambda_{\ell_{1}}^{-}\prod_{p\mid\ell_{1}}\beta_{X^{(\ell_{1},1,1,1)},p}(h)>(1-e^{37-s}(1.083^{10}))\prod_{z_{0}<p\leq z\atop p\mid\prod\alpha_{j}}\beta_{X^{(p,1,1,1)},p}(h).$ (6.12)
One verifies that for $s\geq 38$ we have
$1-e^{37-s}(1.083^{10})>0.$
Thus we have
$\sum_{\ell_{1}\mid P_{h}(z_{0},z)}\Lambda_{\ell_{1}}^{-}\sum_{\ell_{2}\mid P_{h}(z_{0},z)}\lambda_{\ell_{2}}^{+}\sum_{\ell_{3}\mid P_{h}(z_{0},z)}\lambda_{\ell_{3}}^{+}\sum_{\ell_{4}\mid P_{h}(z_{0},z)}\lambda_{\ell_{4}}^{+}\prod_{p\mid\boldsymbol{\ell}}\beta_{X^{\boldsymbol{\ell}},p}(h)\\ \geq(1.083)^{-4}(1-e^{37-s}(1.083^{10}))(\frac{\log(z_{0})}{\log(z)})^{16}+O_{\epsilon}(z_{0}^{\epsilon-1}\log(z)^{16})$
Taking $z_{0}=\log(z)^{33},$ we have
$\sum_{\ell_{1}\mid P_{h}(z_{0},z)}\Lambda_{\ell_{1}}^{-}\sum_{\ell_{2}\mid P_{h}(z_{0},z)}\lambda_{\ell_{2}}^{+}\sum_{\ell_{3}\mid P_{h}(z_{0},z)}\lambda_{\ell_{3}}^{+}\sum_{\ell_{4}\mid P_{h}(z_{0},z)}\lambda_{\ell_{4}}^{+}\prod_{p\mid\boldsymbol{\ell}}\beta_{X^{\boldsymbol{\ell}},p}(h)\gg(\frac{\log(z_{0})}{\log(z)})^{16}.$ (6.13)
Choose suitable $c_{p},$ and again using [6, (3.30)], we conclude that
$W_{\textbf{c},h}(z_{0})\gg\frac{1}{\log(z_{0})^{3}}.$ (6.14)
Thus the main term has a lower bound
$\gg\frac{\log(z_{0})^{13}}{\log(z)^{16}}r_{gen^{+}(X^{\textbf{1}})}(h).$ (6.15)
Thus we conclude
$S_{\textbf{c},h}(A_{\textbf{1}},z)\gg(\frac{\log(z_{0})^{13}}{\log(z)^{16}}+O((\Delta^{-\frac{1}{2}}H(n)^{5}\log(D_{0})^{8}\\ +\Delta^{\epsilon-B}\log(z_{0})^{8}+\Delta^{C}e^{-s_{0}})(\frac{\log(z)}{\log(z_{0})})^{6}))r_{gen^{+}(X^{\boldsymbol{\ell}})}(h)\\ +O_{\epsilon}((DD_{0})^{26+\epsilon}(\prod_{j=1}^{4}\alpha_{j})^{\frac{11}{4}+\epsilon}h^{\frac{1}{2}+\epsilon}.$ (6.16)
Write $z=h^{\theta},s=38,D_{0}=h^{\epsilon},$ so $DD_{0}\ll_{\epsilon}h^{38\theta+\epsilon}$ and $s_{0}\gg_{\epsilon}\frac{\log(h)}{\log(\log(h))}.$ We can choose $\Delta=H(n)^{10}\log(h)^{\max(62,\frac{23}{B})}$. We then note that $\Delta\gg H(n)^{10}\log(h)^{62}$ together with $\log(D_{0})\ll\log(h))$ implies that
$\Delta^{-\frac{1}{2}}H(n)^{5}\log(D_{0})^{8}(\frac{\log(z)}{\log(z_{0})})^{6}\ll_{\epsilon}\log(h)^{-17+\epsilon}\log(z_{0})^{-6}$
while $\Delta\gg\log(h)^{\frac{23}{B}}$ implies
$\Delta^{\epsilon-B}\log(z_{0})^{8}(\frac{\log(z)}{\log(z_{0})})^{6}\ll\log(h)^{-17+\epsilon}\log(z_{0})^{2}$
We can conclude that
$S_{\textbf{c},h}(A_{\textbf{1}},z)\gg\frac{\log(z_{0})^{13}}{\log(z)^{16}}r_{gen^{+}(X^{\boldsymbol{\ell}})}(h)+O_{\epsilon}(h^{988\theta+\frac{1}{2}+\epsilon}(\prod_{j=1}^{4}\alpha_{j})^{\frac{11}{4}+\epsilon})$ (6.17)
Then by Lemma 3.6 we have
$\gg_{\epsilon}(\prod_{j=1}^{4}\alpha_{j})^{-\frac{1}{2}}h^{1-\epsilon}+O(h^{988\theta+\frac{1}{2}+\epsilon}(\prod_{j=1}^{4}\alpha_{j})^{\frac{11}{4}+\epsilon})$ (6.18)
Taking
$\theta<\frac{1}{1977}$
We have
$988\theta+\frac{1}{2}<1,$
and for sufficiently large $h$ we have
$S_{\textbf{c},h}(A_{\textbf{1}},z)>0$
Hence in (1.1) we may take that odd $p\mid x$ implies that
$p>h^{\theta}=(8(m-2)n+\sum_{j=1}^{4}\alpha_{j}(m-4)^{2})^{\theta}\gg_{m}n^{\theta}=n^{\frac{1}{1977}}$
and $\text{ord}_{p}(x_{j})<c_{p}$ for all $1\leq j\leq 4$ If
$\sum_{p>n^{\frac{1}{1900}}}\text{ord}_{p}(x)=a>0,$ (6.19)
then
$p_{m}(x)\geq\frac{(m-2)x^{2}}{2}-\frac{\mid m-4\mid\cdot\mid x\mid}{2}\geq\frac{m-\frac{5}{2}}{2}x^{2},$
Then, letting $\beta$ be the maximum of the respective choices of $a$ for $x_{1},x_{2},x_{3},x_{4}$ from (6.18), we have
$n=\sum_{j=1}^{4}\alpha_{j}p_{m}(x_{j})\gg_{m}n^{\frac{2\beta}{1977}},$
and we conclude that
$\beta\leq 988.$
Overall, we obtain that $x_{1},x_{2},x_{3},x_{4}$ are divisible by at most $988$ primes. ∎
Theorem 6.1.
Let $h=8(m-2)n+\sum_{j=1}^{4}\alpha_{j}(m-4)^{2}$ and $\epsilon>0.$ Then $r_{X}(h)\neq 0$ if
$h\gg_{\epsilon}(2(m-2))^{21+\epsilon}(\prod_{j=1}^{4}\alpha_{j})^{6+\epsilon}.$
Proof.
By Lemma 4.2, we have
$\mid a_{G_{X}}(h)\mid$ $\ll M^{\frac{11}{2}+\epsilon}N_{\boldsymbol{\alpha}}^{\frac{5}{2}+\epsilon}h^{\frac{1}{2}+\epsilon}$
$=(2(m-2))^{\frac{11}{2}}(4(m-2)^{2}\prod_{j=1}^{4}\alpha_{j})^{\frac{5}{2}}h^{\frac{1}{2}+\epsilon}.$
Then by comparing with the main term in (6.18), $r_{X}(h)\neq 0$ implies
$(\prod_{j=1}^{4}\alpha_{j})^{-\frac{1}{2}}h^{1-\epsilon}\gg_{\epsilon}(2(m-2))^{\frac{11}{2}}(4(m-2)^{2}\prod_{j=1}^{4}\alpha_{j})^{\frac{5}{2}}h^{\frac{1}{2}+\epsilon}$
Therefore we have
$h\gg_{\epsilon}(2(m-2))^{21+\epsilon}(\prod_{j=1}^{4}\alpha_{j})^{6+\epsilon}.$

### v2

@@
 It is obvious that if $(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4})$ is too arbitrary, i.e.
-$(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4})=(p,p,p,p).$ Then $n$ that can be represented is restricted.
+$(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4})=(p,p,p,p),$ then $n$ that can be represented is restricted.
 We hence restrict ourselves to the case where $\prod_{j=1}^{4}\alpha_{j}$ is odd and square-free.
-To understand the above sum and the relationship with shifted lattice theory, by completing the square, (1.1) is equivalent to
+To understand the above sum and the relationship with shifted lattice theory, we complete the square so that (1.1) becomes
 $\sum_{j=1}^{4}\alpha_{j}X_{j}^{2}=8(m-2)n+\sum_{j=1}^{4}\alpha_{j}(m-4)^{2},$
@@
 Further suppose $\text{ord}_{p}(\prod_{j=1}^{4}\alpha_{j})\leq 1$ for any odd prime $p.$ Then for $n$ sufficiently large, the equation (1.1) is solvable with each $x_{j}$ containing at most $988$ prime factors.
+Remark.
+We further assume that $m-4\not\equiv 0\pmod{3}$ and $m-4\not\equiv 0\pmod{5}$ due to the following technical issue.
+Suppose we have $m-4\equiv 0\pmod{3}$ or $m-4\equiv 0\pmod{5}$, then $X=2(m-2)x+4-m\not\equiv 0\pmod{p}$ is equivalent to $x\not\equiv 0\pmod{p}$.
+Therefore the integers n that can be represented as (1.1) with $x_{1},x_{2},x_{3},x_{4}\not\equiv 0\pmod{p}$ is restricted: when $p=3,$ only those $n$ with $n\equiv 0\pmod{p}$, when $p=5$ only those $n$ with $n\not\equiv 0\pmod{p}$.
 The paper is organized as follows.
@@
 $\Theta_{X}(z):=\sum_{x\in X}q^{Q(x)}=\sum_{n=0}^{\infty}r_{X}(n)q^{n}.$
-Using [1, Proposition 2.1], a straightforward calculation shows that the theta series of cosets of rank $k$ are modular forms of weight $\frac{k}{2}.$ Therefore is splits into a sum of Eisenstein series $E_{X}$ and a cusp form $G_{X}.$ Let $a_{E_{X}}(n)$ and $a_{G_{X}}(n)$ denote the $n$-th Fourier coefficient of $E_{X}$ and $G_{X}$ for any integer $n\geq 0,$ respectively.
+Using [1, Proposition 2.1], a straightforward calculation shows that the theta series of cosets of rank $k$ are modular forms of weight $\frac{k}{2}.$ Therefore it splits into a sum of Eisenstein series $E_{X}$ and a cusp form $G_{X}.$ Let $a_{E_{X}}(n)$ and $a_{G_{X}}(n)$ denote the $n$-th Fourier coefficient of $E_{X}$ and $G_{X}$ for any integer $n\geq 0,$ respectively.
 We have
@@
 Shimura [7, Theorem 1.5] showed that the $h$-th Fourier coefficient of the Eisenstein series part of $\Theta_{X}$ can be expressed as the product of local densities.
-We refer readers to [8, 9] for more details.
+We refer the readers to [8] and [9] for more details.
 We now provide the statement for the special case when $k=4.$
@@
 Let $\lambda_{\textbf{d}}$ be the characteristic function of $X^{\textbf{d}}.$ In our case $e_{1}$ is the product of all primes dividing $2(m-2)\prod_{j=1}^{4}\alpha_{j}d_{j},$ and $e_{1}=e^{\prime}.$ Therefore we have
-$a_{E_{X}}(h)=\frac{(2\pi)^{2}h}{(16d_{L^{\textbf{d}}})^{\frac{1}{2}}\Gamma(2)L(2,\psi)}\cdot\prod_{p\mid e_{1}}\frac{b_{p}(h,\lambda,0)}{(1-\psi(p)p^{-2})}\cdot\prod_{{p\mid h\atop{p\nmid e_{1}}}}\gamma_{p}(2).$ (3.2)
+$a_{E_{X^{\textbf{d}}}}(h)=\frac{(2\pi)^{2}h}{(16d_{L^{\textbf{d}}})^{\frac{1}{2}}\Gamma(2)L(2,\psi)}\cdot\prod_{p\mid e_{1}}\frac{b_{p}(h,\lambda,0)}{(1-\psi(p)p^{-2})}\cdot\prod_{{p\mid h\atop{p\nmid e_{1}}}}\gamma_{p}(2).$ (3.2)
+Furthermore, $t_{p}$ defined in [7, Section 1.6] is equal to 2 in our case. Therefore for any odd prime $p,$ we have
+$\gamma_{p}(2)=\begin{cases}\frac{1-(-p^{1-s})^{v+1}}{1+p^{1-s}},&\text{if $\psi(p)\neq 0$ and $c_{p}\in\mathbb{Z}_{p}^{\times}$,}\\ \frac{1-(-p^{1-s})^{v+2}+p^{2-s}(1-(-p^{1-s})^{v})}{(1+p^{-s})(1+p^{1-s})},&\text{if $\psi(p)\neq 0$ and $c_{p}\not\in\mathbb{Z}_{p}^{\times}$},\\ 1+\xi_{p}(c_{p}h)p^{(v+\sigma)(1-s)},&\text{if $\psi(p)=0$}.\end{cases}$
+Where the numbers $v,c_{p},\sigma$ and the function $\xi_{p}$ are defined in [7, Section 1.6].
 3.1 Computation of local densities of $X^{\textbf{d}}$.
-In this subsection, we compute the local densities $b_{p}(h,\lambda_{\textbf{d}},0)$ of $X^{\textbf{d}}$ in order to evaluate $a_{E_{X^{\textbf{d}}}}(h)$ using (3.2).
-We first compute the case for $p=2.$
+In this subsection, we compute the local densities $b_{p}(h,\lambda_{\textbf{d}},0)$ of $X^{\textbf{d}}$ in order to evaluate $a_{E_{X^{\textbf{d}}}}(h)$ using (3.2). We first require a formula for
+$\tau_{\alpha_{j},p,s}(\sigma):=\int_{\mathbb{Z_{p}}}e_{p}(4(m-2)s\sigma\alpha_{j}((m-2)sx^{2}-(m-4)x))\,dx$
+for $\sigma\in\mathbb{Q}_{p}.$ Then by [1, Lemma 3.2], we have
+Lemma 3.2.
+If $\alpha_{j}s\in\mathbb{Z}_{p}$ satisfies $p\mid 2(m-2)\alpha_{j}s,$ then we have
+$\tau_{\alpha_{j},p,s}(\sigma)=\begin{cases}1,&\text{if $4(m-2)\alpha_{j}s\sigma\in\mathbb{Z}_{p}$,}\\ 0,&\text{otherwise.}\end{cases}$
+We first compute the local density for the case $p=2.$
 Lemma 3.3.
 The local density if $X^{\textbf{d}}$ at the prime 2 is given by
-$b_{p}(h,\lambda_{\textbf{d}},0)=\begin{cases}2^{2+\text{min}_{j}\{\text{ord}_{2}(\alpha_{j}d_{j})\}},&\text{if $8(m-2)n\in 4\gcd(\alpha_{1}d_{1},\alpha_{2}d_{2},\alpha_{3}d_{3},\alpha_{4}d_{4})\mathbb{Z}_{2}$,}\\ 0,&\text{otherwise.}\end{cases}$
+$b_{2}(h,\lambda_{\textbf{d}},0)=\begin{cases}2^{2+\text{min}_{j}\{\text{ord}_{2}(\alpha_{j}d_{j})\}},&\text{if $8(m-2)n\in 4\gcd(\alpha_{1}d_{1},\alpha_{2}d_{2},\alpha_{3}d_{3},\alpha_{4}d_{4})\mathbb{Z}_{2}$,}\\ 0,&\text{otherwise.}\end{cases}$
 Proof.
-By [1 Lemma 3.3]. ∎
-For odd $p\mid(m-2)$
+Note that we have
+$Q(u+\nu)=\sum_{j=1}^{4}4\alpha_{j}(m-2)^{2}(d_{j}x_{j}-\frac{m-4}{2(m-2)})^{2},$ (3.3)
+where $\nu=x_{1}(d_{1}e_{1})+x_{2}(d_{2}e_{2})+x_{3}(d_{3}e_{3})+x_{4}(d_{4}e_{4}).$ Then by (3.1) and Lemma 3.2, we have
+$b_{2}(h,\lambda_{\textbf{d}},0)$ $=\int_{\mathbb{Q}_{p}}e_{p}(\sigma(\sum_{j=1}^{4}\alpha_{j}(m-4)^{2}-h))\prod_{j=1}^{4}\tau_{\alpha_{j},2,d_{j}}(\sigma)\,d\sigma$
+$=\int_{\frac{1}{4\gcd(\alpha_{1}d_{1},\alpha_{2}d_{2},\alpha_{3}d_{3},\alpha_{4}d_{4})}\mathbb{Z}_{2}}e_{p}(-8n(m-2)\sigma)\,d\sigma$
+$=\begin{cases}2^{2+\text{min}_{j}\{\text{ord}_{2}(\alpha_{j}d_{j})\}},&\text{if $8(m-2)n\in 4\gcd(\alpha_{1}d_{1},\alpha_{2}d_{2},\alpha_{3}d_{3},\alpha_{4}d_{4})\mathbb{Z}_{2}$,}\\ 0,&\text{otherwise.}\end{cases}$
+We now compute the local density for odd $p\mid(m-2).$
 Lemma 3.4.
@@
 Proof.
-By [1 Lemma 3.4]. ∎
+By (3.1), Lemma 3.2 and (3.3), we have
+$b_{p}(h,\lambda_{\textbf{d}},0)$ $=\int_{\mathbb{Q}_{p}}e_{p}(\sigma(\sum_{j=1}^{4}\alpha_{j}(m-4)^{2}-h))\prod_{j=1}^{4}\tau_{\alpha_{j},2,d_{j}}(\sigma)\,d\sigma$
+$=\int_{\frac{1}{(m-2)\gcd(\alpha_{1}d_{1},\alpha_{2}d_{2},\alpha_{3}d_{3},\alpha_{4}d_{4})}\mathbb{Z}_{p}}e_{p}(-8n(m-2)\sigma)\,d\sigma$
+$=\begin{cases}p^{\text{ord}_{p}(m-2)+\text{min}_{j}\{\text{ord}_{p}(\alpha_{j}d_{j})\}},&\text{if $8(m-2)n\in\gcd(\alpha_{1}d_{1},\alpha_{2}d_{2},\alpha_{3}d_{3},\alpha_{4}d_{4})\mathbb{Z}_{p}$,}\\ 0,&\text{otherwise.}\end{cases}$
 For $p\geq 3,$ we will use the formula from [5, Theorem 4.2] to compute the local densities. We provide the statement for $r=4$ in the following.
@@
 Then we have
-$I_{p}(n;\phi)=1+(1-\frac{1}{p})\sum_{1\leq t\leq\text{min}(\textbf{t}_{d},\textbf{t}_{n})\atop l_{p}(t)\text{ even}}\delta_{p}(t)p^{\tau_{p}(t)}+\delta_{p}(\textbf{t}_{n}+1)w_{p}p^{\tau_{p}(\textbf{t}_{n}+1)},$
+$b_{p}(h,\lambda_{\textbf{d}},0)=1+(1-\frac{1}{p})\sum_{1\leq t\leq\text{min}(\textbf{t}_{d},\textbf{t}_{n})\atop l_{p}(t)\text{ even}}\delta_{p}(t)p^{\tau_{p}(t)}+\delta_{p}(\textbf{t}_{n}+1)w_{p}p^{\tau_{p}(\textbf{t}_{n}+1)},$
 where for any integer $t\in\mathbb{Z},$ we define
@@
 Lemma 3.6.
-Let $p$ be an odd prime such that $p\nmid(m-2)(m-4),p\mid\prod_{j=1}^{4}d_{j},$ and $p\nmid\prod_{j=1}^{4}\alpha_{j}$
+Let $p$ be an odd prime such that $p\nmid(m-2)(m-4),p\mid\prod_{j=1}^{4}d_{j},$ and $p\nmid\prod_{j=1}^{4}\alpha_{j}.$
 When $\mid N_{p}\mid=0,$ we have
@@
 $1-\frac{1}{p}\leq b_{p}(h,\lambda_{\textbf{d}},0)\leq 1+\frac{1}{p}.$
-Lower bound for the $\textbf{d}=1$ case. We are now ready to give a lower bound for $a_{E_{X^{\textbf{1}}}(h)}.$
+Proof.
+For $\mid N_{p}\mid=0,$ we have $\textbf{t}_{d}=1$ and $L_{p}(t)=\emptyset,$ thus $\delta_{p}(t)=1$ and $\tau_{p}(t)=t.$ Therefore
+$b_{p}(h,\lambda_{\textbf{d}},0)$ $=1+(1-\frac{1}{p})\sum_{1\leq t\leq\text{min}(\textbf{t}_{d},\textbf{t}_{n})}p^{t}+w_{p}\cdot p^{\textbf{t}_{n}+1}$
+$=\begin{cases}p&\text{if $N\in p\mathbb{Z}_{p}$,}\\ 0&\text{if $N\not\in p\mathbb{Z}_{p}$.}\end{cases}$
+For $\mid N_{p}\mid=1,$ wlog we assume $1\in N_{p}.$ Then $t_{1}=0.$ Now the computation splits into two cases, which is $t_{1}<\text{min}(\textbf{t}_{d},\textbf{t}_{n})$ or $t_{1}=\text{min}(\textbf{t}_{d},\textbf{t}_{n}).$ For the first case, as we assume $0=t_{1}<\text{min}(\textbf{t}_{d},\textbf{t}_{n}),$ thus $1=\textbf{t}_{d}\leq\textbf{t}_{n},$ hence $\text{min}(\textbf{t}_{d},\textbf{t}_{n})=1$ and $w_{p}=0$.
+Now as $l_{p}(1)=1,$ the sum in the middle is empty and the local density is
+$b_{p}(h,\lambda_{\textbf{d}},0)=1$
+For $t_{1}=\text{min}(\textbf{t}_{d},\textbf{t}_{n}),$ we have $\textbf{t}_{d}>\textbf{t}_{n}=0.$ Therefore the local density is bounded by
+$0\leq b_{p}(h,\lambda_{\textbf{d}},0)=1+(\frac{u_{1}}{p})(\frac{\textbf{u}_{n}}{p})\leq 2.$
+The remaining cases can be compute similarly. ∎
+We are now ready to give a lower bound for $a_{E_{X^{\textbf{1}}}(h)}.$
 Lemma 3.7.
@@
 $\begin{split}&\frac{a_{E_{X^{\boldsymbol{d\ell}}}(h)}}{a_{E_{X^{\boldsymbol{\ell}}}(h)}}\\ &=\frac{1}{d_{1}d_{2}d_{3}d_{4}}\cdot\prod_{p\mid(m-2)}\frac{b_{p}(h,\lambda_{\boldsymbol{d\ell}},0)}{b_{p}(h,\lambda_{\boldsymbol{\ell}},0)}\cdot\prod_{p\mid d_{1}d_{2}d_{3}d_{4}\atop p\nmid 2(m-2)}\frac{b_{p}(h,\lambda_{\textbf{d}},0)}{1-\psi(p)p^{-2}}\cdot\prod_{p\mid(h,d_{1}d_{2}d_{3}d_{4})\atop p\nmid 2(m-2)}\gamma_{p}^{-1}(2).\end{split}$ (3.4)
+Remark.
+Note that (3.4) follows from the fact that for general $\textbf{d}\in\mathbb{N}^{4},$ we have $b_{p}(h,\lambda_{\textbf{d}},0)=b_{p}(h,\lambda_{p^{\text{ord}_{p}\textbf{d}}},0).$
 Define
-$\begin{split}\beta_{X^{\textbf{p}^{c}},p}(h):=\frac{1}{p^{c_{1}+c_{2}+c_{3}+c_{4}}}&(\frac{b_{p}(h,\lambda_{\textbf{p}^{c}},0)}{b_{p}(h,\lambda_{\textbf{1}},0)})^{\delta_{p\mid(m-2)}}\\ &\cdot(b_{p}(h,\lambda_{\textbf{p}^{c}},0)\cdot\frac{1}{1-\psi(p)p^{-2}}\cdot\gamma_{p}^{-1}(2))^{\delta_{p\nmid 2(m-2),c\neq 0}}\end{split}$ (3.4)
-Therefore we have
+$\begin{split}\beta_{X^{\textbf{p}^{\textbf{c}}},p}(h):=\frac{1}{p^{c_{1}+c_{2}+c_{3}+c_{4}}}&(\frac{b_{p}(h,\lambda_{\textbf{p}^{c}},0)}{b_{p}(h,\lambda_{\textbf{1}},0)})^{\delta_{p\mid(m-2)}}\\ &\cdot(b_{p}(h,\lambda_{\textbf{p}^{\textbf{c}}},0)\cdot\frac{1}{1-\psi(p)p^{-2}}\cdot\gamma_{p}^{-1}(2))^{\delta_{p\nmid 2(m-2),\textbf{c}\neq\textbf{0}}}\end{split}$ (3.5)
+where $p^{\textbf{c}}=(p^{c_{1}},p^{c_{2}},p^{c_{3}},p^{c_{4}})$ for $\textbf{c}=(c_{1},c_{2},c_{3},c_{4}).$ Therefore we have
 $\frac{a_{E_{X^{\boldsymbol{d\ell}}}(h)}}{a_{E_{X^{\boldsymbol{\ell}}}(h)}}=\prod_{p\text{ odd}}\beta_{X^{\textbf{p}^{ord_{p}(\textbf{d})}},p}(h),$ (3.6)
@@
 $\beta_{X^{(p,1,1,1),p}}(h)$ $=\frac{1}{p}w_{p}(h)b_{p}(h,\lambda_{(p,1,1,1)},0)$
-$\leq\frac{2}{p}$
-and similarly for other cases. ∎
+$\leq\frac{2}{p}.$
+Other cases can be bound similarly. ∎
 Lemma 3.9.
@@
 $\max\{\beta_{X^{(p,p,1,1)},p}(h),\beta_{X^{(p,1,p,1)},p}(h),\beta_{X^{(p,1,1,p)},p}(h),\beta_{X^{(1,p,p,1)},p}(h)\\ ,\beta_{X^{(1,p,1,p)},p}(h),\beta_{X^{(1,1,p,p)},p}(h)\}\leq\frac{2p}{(p-1)^{2}(p+1)}\leq\frac{16}{p^{2}}$
-$\max\{\beta_{X^{(p,p,p,1)},p}(h),\beta_{X^{(p,p,1,p)},p}(h),\beta_{X^{(p,1,p,p)},p}(h),\beta_{X^{(1,p,p,p)},p}(h)\}\leq\frac{2}{(p-1)^{2}(p+1)}\leq\frac{64}{p^{3}}$
+$\max\{\beta_{X^{(p,p,p,1)},p}(h),\beta_{X^{(p,p,1,p)},p}(h),\beta_{X^{(p,1,p,p)},p}(h)\\ ,\beta_{X^{(1,p,p,p)},p}(h)\}\leq\frac{2}{(p-1)^{2}(p+1)}\leq\frac{64}{p^{3}}$
 $\beta_{X^{(p,p,p,p)},p}(h)\leq\begin{cases}\frac{1}{(p-1)^{2}(p+1)},&\text{for $n\in p\mathbb{Z}_{p}$ }\\ \frac{256}{p^{4}},&\text{for $n\not\in p\mathbb{Z}_{p}$ }\end{cases}$
 Proof.
-By Theorem 3.4, we have
+By Theorem 3.5, the maximum value of $b_{p}(h,\lambda_{p^{\textbf{c}}},0)$ where $\textbf{c}\in\{(0,0,0,1),(0,0,1,0),(0,1,0,0),(1,0,0,0)\}$ is achieved when $\textbf{t}_{d}=1,t_{1}=t_{2}=0,t_{3}=1,$ where $N_{p}=\{1,2,3\},$ and $\textbf{t}_{n}\geq\textbf{t}_{d}.$ Then
+$b_{p}(h,\lambda_{p^{\textbf{c}}},0)=1+(1-\frac{1}{p})\delta_{p}(1)\leq 2-\frac{1}{p}.$
+Therefore we have
 $\max\{\beta_{X^{(p,1,1,1)},p}(h),\beta_{X^{(1,p,1,1)},p}(h),\beta_{X^{(1,1,p,1)},p}(h),\beta_{X^{(1,1,1,p)},p}(h)\}\\ \leq\frac{1}{p}\cdot(2-\frac{1}{p})\cdot\frac{p^{3}}{(p-1)^{2}(p+1)}\leq\frac{4}{p}\text{ for }p\geq 5.$
-and similarly for other cases. ∎
+For $\textbf{c}\in\{(0,0,1,1),(0,1,0,1),(1,0,0,1),(0,1,1,0),(1,0,1,0),(1,1,0,0)\},$ the maximum value of $b_{p}(h,\lambda_{p^{\textbf{c}}},0)$ is achieved when $\textbf{t}_{d}=1,t_{1}=0,t_{2}=1,$ where $N_{p}=\{1,2\},$ $t_{3}=t_{4}=1,$ where $D_{p}=\{3,4\},$ and $\textbf{t}_{n}<\textbf{t}_{d}.$ Then
+$b_{p}(h,\lambda_{p^{\textbf{c}}},0)=1+(\frac{u_{1}}{p})(\frac{\textbf{u}_{n}}{p})\leq 2.$
+Therefore we have
+$\max\{\beta_{X^{(p,p,1,1)},p}(h),\beta_{X^{(p,1,p,1)},p}(h),\beta_{X^{(p,1,1,p)},p}(h),\\ \beta_{X^{(1,p,p,1)},p}(h),\beta_{X^{(1,p,1,p)},p}(h),\beta_{X^{(1,1,p,p)},p}(h)\}\\ \leq\frac{1}{p^{2}}\cdot 2\cdot\frac{p^{3}}{(p-1)^{2}(p+1)}\leq\frac{16}{p^{2}}.$
+For $\textbf{c}\in\{(0,1,1,1),(1,0,1,1),(1,1,0,1),(1,1,1,0)\},$ the maximum value of $b_{p}(h,\lambda_{p^{\textbf{c}}},0)$ is achieved when $\textbf{t}_{d}=1,t_{1}=0$ where $N_{p}=\{1\},$ and $\textbf{t}_{n}<\textbf{t}_{d}.$ Then
+$b_{p}(h,\lambda_{p^{\textbf{c}}},0)=1+(\frac{u_{1}}{p})(\frac{\textbf{u}_{n}}{p})\leq 2.$
+Therefore we have
+$\max\{\beta_{X^{(p,p,p,1)},p}(h),\beta_{X^{(p,p,1,p)},p}(h),\beta_{X^{(p,1,p,p)},p}(h)\\ ,\beta_{X^{(1,p,p,p)},p}(h)\}\leq\frac{1}{p^{3}}\cdot 2\cdot\frac{p^{3}}{(p-1)^{2}(p+1)}\leq\frac{64}{p^{3}}$
+And finally for $\textbf{c}=(1,1,1,1),$ we have
+$b_{p}(h,\lambda_{p^{\textbf{c}}},0)=\begin{cases}p,&\text{for $n\in p\mathbb{Z}_{p}$}\\ 0,&\text{for $n\not\in p\mathbb{Z}_{p}$}\end{cases}$
+Therefore we have
+$\beta_{X^{(p,p,p,p)},p}(h)\leq\begin{cases}\frac{1}{(p-1)^{2}(p+1)},&\text{for $n\in p\mathbb{Z}_{p}$ }\\ \frac{256}{p^{4}},&\text{for $n\not\in p\mathbb{Z}_{p}$ }\end{cases}$
 4 Error from Cusp form
@@
 $\mid a_{G_{X}}(n)\mid\leq\frac{54}{\pi^{2}\delta_{M^{2}N_{\boldsymbol{\alpha}}}^{\frac{3}{2}}}\frac{M^{2}N_{\boldsymbol{\alpha}}^{2+2\delta}\sqrt{\frac{2\pi}{3}}e^{2\pi}\zeta(1+4\delta)^{\frac{1}{2}}c_{\delta}^{\frac{5}{2}}\varphi(M)}{\prod_{p\mid M,p\nmid N_{\boldsymbol{\alpha}}}(1-p^{-2})^{\frac{1}{2}}\prod_{p\mid N_{\boldsymbol{\alpha}}}(1-p^{-1})^{\frac{1}{2}}}C_{\epsilon}n^{\frac{1}{2}+\epsilon}\\ \times(\sum_{d\mid M^{2}N_{\boldsymbol{\alpha}}}\varphi(\frac{M^{2}N_{\boldsymbol{\alpha}}}{d})\varphi(d)\frac{M^{2}N_{\boldsymbol{\alpha}}}{d}(\frac{\gcd(M^{2},d)}{M^{2}})^{4})^{\frac{1}{2}}(\frac{27}{\Delta_{\boldsymbol{\alpha}}}\frac{M^{2}N_{\boldsymbol{\alpha}}}{\pi\delta_{M^{2}N_{\boldsymbol{\alpha}}}}+16)^{\frac{1}{2}}.$
-Using Lemma 4.1, we have the following bound.
+We next bound the cuspidal contribution.
 Lemma 4.2.
+Using Lemma 4.1, we have
 $\mid a_{G_{X}}(n)\mid\ll M^{\frac{11}{2}+\epsilon}N_{\boldsymbol{\alpha}}^{\frac{5}{2}+\epsilon}n^{\frac{1}{2}+\epsilon}.$